### v3

# Endomorphism algebras of abelian varieties with large cyclic 2-torsion field over a given field

Abstract.
In this article we study the endomorphism algebras of abelian varieties $A$ defined over a given number field $K$ with large cyclic 2-torsion fields. A key step in doing so is to provide criteria for all the endomorphisms of $A$ to be defined over $K(A[2])$, the field generated by its 2-torsion.
When $K=\mathbb{Q}$ and $\operatorname{{\rm Gal}}(\mathbb{Q}(A[2])/\mathbb{Q})$ is cyclic of prime order $p=2\dim(A)+1$, we prove the somewhat surprising result that there are only finitely many possibilities for the geometric endomorphism algebra $\operatorname{{\rm End}}(A)\otimes\mathbb{Q}$. In fact, when $\dim(A)\not\in\{3,5,9,21,33,81\}$, we show $\operatorname{{\rm End}}(A)\otimes\mathbb{Q}$ is a proper subfield of the $p$-th cyclotomic field.
In particular, when $g=2$, $\operatorname{{\rm End}}(A)\otimes\mathbb{Q}$ is isomorphic to either $\mathbb{Q}$ or $\mathbb{Q}(\sqrt{5})$.
1. Introduction
Faltings’ Isogeny Theorem tells us that for an abelian variety $A$, defined over a number field $K$, we can recover its ring of endomorphisms defined over $K$, $\operatorname{{\rm End}}_{K}(A)$, tensored by $\mathbb{Z}_{\ell}$ (where $\ell$ is a prime) from looking at the action of the absolute Galois group $G_{K}\coloneqq\operatorname{{\rm Gal}}(\bar{K}/K)$ on its $\ell$-adic Tate module $T_{\ell}(A)$. More precisely, Faltings’ Isogeny Theorem tells us that the natural injection
$\operatorname{{\rm End}}_{K}(A)\otimes\mathbb{Z}_{\ell}\hookrightarrow\operatorname{{\rm End}}_{K}(T_{\ell}(A))$
is an isomorphism.
The mod $\ell$ version of the above is known to fail to be an isomorphism in general, that is, generally speaking, the injection
$\operatorname{{\rm End}}_{K}(A)\otimes\mathbb{Z}/\ell\mathbb{Z}\hookrightarrow\operatorname{{\rm End}}_{K}(A[\ell])$
is not an isomorphism. However, when the action of $G_{K}$ on $A[\ell]$ is particularly ‘large’ one might hope that $\operatorname{{\rm End}}_{K}(A[\ell])\;(=\operatorname{{\rm End}}_{\bar{K}}(A[\ell])^{G_{K}})$ is small enough to place significant restrictions on $\operatorname{{\rm End}}_{K}(A)\otimes\mathbb{Z}/\ell\mathbb{Z}$ and hence $\operatorname{{\rm End}}_{K}(A)$.
Zarhin has had much success with carrying this out for jacobians of hyperelliptic curves (with $\ell=2$), see for example [Zar00, Zar01, Zar05]. The attraction of this case is that for a hyperelliptic curve determined by a smooth affine model $y^{2}=f(x)$, the Galois group $\operatorname{{\rm Gal}}(K(A[2])/K)$ is equal to $\operatorname{{\rm Gal}}(K(f)/K)$, the Galois group of the splitting field of $f$, and its action on $A[2]$ can be described explicitly via its action on the roots of $f$. Thus giving explicit examples.
In the cases considered by Zarhin, all of the Galois groups are insoluble and act 2-transitively on the roots of $f$. In [Goo21] the author considered restrictions on $\operatorname{{\rm End}}_{K}(A)$ and $\operatorname{{\rm End}}(A)\coloneqq\operatorname{{\rm End}}_{\bar{K}}(A)$ when $\operatorname{{\rm Gal}}(K(A[\ell])/K)$ contained merely an element of large prime order, with no restrictions on $K$.
However, considering the case of elliptic curves, for example, where the possible endomorphism algebras depend solely on $[K:\mathbb{Q}]$, it is natural to impose conditions on $K$. The main advantage in doing so is that this allows for more arithmetic input, whereas the proofs of the above results are heavily representation theoretic.
Thus, in this paper we study restrictions on $\operatorname{{\rm End}}(A)$ (and $\operatorname{{\rm End}}_{K}(A)$) coming from the ground field in conjunction with the results of [Goo21]. In this way, this paper may be seen as a natural continuation of [Goo21].
As an example, we obtain the somewhat surprising [^1] At least in the author’s opinion. result (where $\zeta_{n}$ denotes a primitive $n$-th root of unity):
Theorem 1.1 ($\subset$ Corollary 2.6 + Theorem 2.10).
Let $A/K$ be an abelian variety of dimension $g\geq 2$ with $p=2g+1$ prime. Suppose $\operatorname{{\rm Gal}}(K(A[2])/K)$ has order $p$ and either $K=\mathbb{Q}$ or $K$ is an imaginary quadratic field such that $p\nmid\#\mathrm{Cl}(K)$. Then either
  • $\operatorname{{\rm End}}(A)\otimes\mathbb{Q}$ is a proper subfield of $\mathbb{Q}(\zeta_{p})$; or
  • $p\in\{7,11,19,43,67,163\}$ and $A$ is isogenous over $\bar{K}$ to the power of an elliptic curve with complex multiplication by $\mathbb{Q}(\sqrt{-p})$.
In particular there are only finitely many possibilities for $\operatorname{{\rm End}}^{0}(A)$.
What is particularly interesting here is that this does not fit into the philosophy of the above results, which tell us to expect that the bigger the Galois group $\operatorname{{\rm Gal}}(K(A[2])/K)$ is, the smaller the endomorphism algebra of $A$ should be. Indeed, the jacobian defined by $y^{2}=x^{5}-2$ has CM by $\mathbb{Q}(\zeta_{5})$ and the Galois group of the splitting field of $x^{5}-2$ has order 20. Whereas by the above, an abelian surface $A$ over $\mathbb{Q}$ with $\operatorname{{\rm Gal}}(\mathbb{Q}(A[2])/\mathbb{Q})$ of order 5 must have endomorphism algebra a proper subfield of $\mathbb{Q}(\zeta_{5})$. This is of course only possibly because we are dealing the 2-torsion opposed to the entire 2-adic Tate module.
Intuitively, the reason for this discrepancy is already exemplified in the above example: $\operatorname{{\rm Gal}}(K(A[2])/K)$ being small should force the endomorphisms to be defined over a small extension of $K$. To make this rigorous, we need to study the minimal extension $L/K$ over which $A$ obtains all its endomorphisms. This field was christened in [GK17] as the endomorphism field of $A$
An important result in understanding this field was obtained by Silverberg, who showed it is contained in $K(A[n])$ for $n\geq 3$ [Sil92, Theorem 2.4]. Bounds on $[L\colon K]$ were then studied in [GK17] and also in [FKRS12, FKS21] via its connection to Sato-Tate groups.
However, as eluded to above, we wish to be able to bound $L$ relative to $K(A[2])$. The following theorem allows us to do so (see also Theorems 3.3, 3.5 for an analogue in the setting of QM surfaces):
Theorem 1.2 ($=$ Theorem 2.1).
Suppose $E=\operatorname{{\rm End}}(A)\otimes\mathbb{Q}$ is a (finite) Galois extension of $\mathbb{Q}$ and $L\nsubseteq K(A[2])$. The following hold:
  • $\operatorname{{\rm Gal}}(E/\mathbb{Q})$ has a non-trivial normal elementary abelian 2-subgroup;
  • if $\operatorname{{\rm End}}(A)$ is 2-maximal in $E$, then $2$ is wildly ramified in $E/\mathbb{Q}$.
In particular, if $E/\mathbb{Q}$ is Galois, $\operatorname{{\rm End}}(A)$ is a 2-maximal order and $2$ is not wildly ramified, then $L\subseteq K(A[2])$.
Recall that an order $\mathcal{O}\subseteq\operatorname{{\rm End}}(A)\otimes\mathbb{Q}$ is said to be 2-maximal if for any other order $\mathcal{O}^{\prime}\subseteq\operatorname{{\rm End}}(A)\otimes\mathbb{Q}$ satisfying $\mathcal{O}\otimes\mathbb{Z}_{2}\subseteq\mathcal{O}^{\prime}\otimes\mathbb{Z}_{2}$ we have $\mathcal{O}\otimes\mathbb{Z}_{2}=\mathcal{O}^{\prime}\otimes\mathbb{Z}_{2}$.
Example 2.2 shows that the condition $\operatorname{{\rm End}}(A)$ is 2-maximal cannot be removed.
Finally, we would like to end the introduction with a challenge [^2] See also Question 2.12 for the reader which we hope will provide some impetus for studying moduli spaces of abelian surfaces. To do so let us state the following corollary of Theorems 1.1,1.2:
Corollary 1.3 ($=$ Corollary 2.8).
Let $A/\mathbb{Q}$ be an abelian surface. Suppose $\operatorname{{\rm Gal}}(\mathbb{Q}(A[2])/\mathbb{Q})\cong C_{5}$. Then either $\operatorname{{\rm End}}(A)=\mathbb{Z}$ or $\operatorname{{\rm End}}_{\mathbb{Q}}^{0}(A)=\operatorname{{\rm End}}^{0}(A)=\mathbb{Q}(\sqrt{5})$.
Both cases are possible, see Example 2.9.
Challenge 1.4.
Give an independent proof of Corollary 1.3 using moduli spaces of abelian surfaces.
Acknowledgements.
The author thanks Tim Dokchitser, Francesc Fité, Xavier Guitart, Jef Laga, Jeremy Rickard, Víctor Rotger, Ciaran Schembri, John Voight and David Zywina for useful comments, questions and discussions.
The author thanks the University of Bristol and the Université Clermont Auvergne for their hospitality during the period this paper was written. Where the author was funded by the EPSRC (grant number EP/N509619/1) and an early career fellowship from the LMS respectively. Finally, the author thanks Universitat de Barcelona where final revisions to this paper were made.
Notation.
Throughout $A$ will denote an abelian variety of dimension $g$ defined over a number field $K$ (often $\mathbb{Q}$). Moreover, we denote by $L$ the smallest extension of $K$ over which all the endomorphisms of $A$ are defined.
We write $\operatorname{{\rm End}}_{F}(A)$ for the ring of endomorphisms of $A$ defined over $F/K$, $\operatorname{{\rm End}}(A)\coloneqq\operatorname{{\rm End}}_{\bar{K}}(A)$ and $\operatorname{{\rm End}}_{F}^{0}(A)\coloneqq\operatorname{{\rm End}}_{F}(A)\otimes\mathbb{Q}$.
We denote by $C_{n},D_{n},F_{n}$ the cyclic group of order $n$, the dihedral group of order $2n$, and the Frobenius group isomorphic to $\mathbb{F}_{n}\rtimes\mathbb{F}_{n}^{*}$ (for $n$ a prime power) respectively.
A primitive $n$-th root of unity is denoted by $\zeta_{n}$. For a polynomial $f\in K[x]$, we denote its splitting field by $K(f)$, and write $\operatorname{{\rm Gal}}(f)$ for $\operatorname{{\rm Gal}}(K(f)/K)$.
2. Proof of the main theorems
Silverberg has shown the endomorphism field $L/K$ of an abelian variety $A/K$ is contained in the $n$-torsion field for every $n\geq 3$ [Sil92, Theorem 2.4]. The case $n=4$ implies $\operatorname{{\rm Gal}}(L/L\cap K(A[2]))$ is an elementary abelian 2-group, since $\operatorname{{\rm Gal}}(K(A[4])/K(A[2]))$ is too, for details see [Goo21, Proposition 3.9]. Below we provide conditions on $\operatorname{{\rm End}}(A)$ which guarantee $L\subseteq K(A[2])$.
Theorem 2.1.
Suppose $E=\operatorname{{\rm End}}^{0}(A)$ is a (finite) Galois extension of $\mathbb{Q}$ and $L\nsubseteq K(A[2])$. The following hold:
  • $\operatorname{{\rm Gal}}(E/\mathbb{Q})$ has a non-trivial normal elementary abelian 2-subgroup;
  • if $\operatorname{{\rm End}}(A)$ is 2-maximal in $E$, then $2$ is wildly ramified in $E/\mathbb{Q}$.
In particular, if $E/\mathbb{Q}$ is Galois, $\operatorname{{\rm End}}(A)$ is a 2-maximal order and $2$ is not wildly ramified, then $L\subseteq K(A[2])$.
Proof.
If $\operatorname{{\rm Gal}}(L/K)$ acts faithfully on $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}/2\mathbb{Z}$, then we may recuperate Lemma 2.1 in [Sil92] with $N=2$, $\Lambda=\operatorname{{\rm End}}(A)$ and $G$ the image of $\operatorname{{\rm Gal}}(L/K)$ in $\operatorname{{\rm Aut}}(\Lambda)$. Following through the proofs of Propositions 2.2, 2.3 and Theorem 2.4 in [Sil92] shows $L\subseteq K(A[2])$. Hence we may suppose the action of $\operatorname{{\rm Gal}}(L/K)$ on $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}/2\mathbb{Z}$ has a non-trivial element $\gamma$ in its kernel.
Since $\operatorname{{\rm Gal}}(L/K)$ acts on $\operatorname{{\rm End}}(A)$ by field automorphisms of $\operatorname{{\rm End}}^{0}(A)$, we may view its image as a subgroup of $\operatorname{{\rm Gal}}(E/\mathbb{Q})$. Identify $\gamma$ with its image in $\operatorname{{\rm Gal}}(E/\mathbb{Q})$. Let $d$ be the rank of $\operatorname{{\rm End}}(A)$ as a free $\mathbb{Z}$-module and view $\operatorname{{\rm Gal}}(E/\mathbb{Q})$ as a (finite) subgroup of $\operatorname{{\rm GL}}_{d}(\mathbb{Z})$ via its action on $\operatorname{{\rm End}}(A)$. The reduction map $\operatorname{{\rm GL}}_{d}(\mathbb{Z})\rightarrow\operatorname{{\rm GL}}_{d}(\mathbb{Z}/4\mathbb{Z})$ is injective on finite groups and the kernel of $\operatorname{{\rm GL}}_{d}(\mathbb{Z}/4\mathbb{Z})\rightarrow\operatorname{{\rm GL}}_{d}(\mathbb{Z}/2\mathbb{Z})$ is an elementary abelian $2$-group (for more details see [Goo21, Prop. 3.9]). Thus the kernel of the map $\operatorname{{\rm Gal}}(E/\mathbb{Q})\rightarrow\operatorname{{\rm Aut}}(\operatorname{{\rm End}}(A)\otimes\mathbb{Z}/2\mathbb{Z})\cong\operatorname{{\rm GL}}_{d}(\mathbb{Z}/2\mathbb{Z})$ is an elementary abelian $2$-group which is non-trivial since it contains $\gamma$.
Assume now $\operatorname{{\rm End}}(A)$ is a 2-maximal order in $E$. As $\gamma$ acts trivially on the order $\operatorname{{\rm End}}(A)$ modulo $2$, it also acts trivially on $\operatorname{{\rm End}}(A)$ modulo every prime above $2$ in $E$. Since $\operatorname{{\rm End}}(A)$ is 2-maximal, we find $\gamma$ belongs to the inertia group $I_{\mathfrak{p}}$ for every prime $\mathfrak{p}$ above 2 in $E$. By the above $\gamma$ has order two, and hence $2$ is wildly ramified in $E/\mathbb{Q}$. ∎
Example 2.2.
The condition that $\operatorname{{\rm End}}(A)$ is 2-maximal cannot be removed. Indeed, the elliptic curve $y^{2}=(x+2)(x^{2}-2x-11)$ has CM by $\mathbb{Z}[\sqrt{-3}]$ and its 2-torsion field is $\mathbb{Q}(\sqrt{3})$, see [Sil94, Appendix A §3].
Proposition 2.3.
Let $A/K$ be an abelian variety of dimension $g\geq 1$ with $p=2g+1$ prime. Suppose $\mathfrak{q}$ is a prime of bad reduction for $A$ and there is an element of order $p$ in the image of the inertia group $I_{\mathfrak{q}}$ on $A[\ell]$ for some $\mathfrak{q}\nmid\ell$. Then either
  • $p$ does not divide $[L:K]$ and $\operatorname{{\rm End}}^{0}(A)$ is a subfield of $\mathbb{Q}(\zeta_{p})$; or
  • $p$ divides $[L:K]$, $g\geq 3$ and $A$ is isogenous over $\bar{K}$ to the power of an absolutely simple abelian variety with complex multiplication by a proper subfield of $\mathbb{Q}(\zeta_{p})$.
In particular there are only finitely many possibilities for $\operatorname{{\rm End}}^{0}(A)$.
Proof.
Suppose $p$ divides $[L:\mathbb{Q}]$. Then by Theorem 2.5 of [Goo21], $g\geq 3$ and $A$ is isogenous over $\bar{K}$ to the power of an absolutely simple abelian variety with complex multiplication by a proper subfield $F$ of $\mathbb{Q}(\zeta_{p})$. Hence we may suppose $p$ does not divide $[L:K]$.
The criterion of Néron-Ogg-Shafarevich implies that the image of the inertia group $I_{\mathfrak{q}}$ on the Tate module $T_{\ell}(A)$ for any prime $\ell$, not divisible by $\mathfrak{q}$, contains an element $\tau$ of order $p$. Furthermore, as the trace of $\tau$ is an integer [ST68, Thm. 2] its eigenvalues are the primitive $p$-th roots of unity (owing to the fact $\tau$ has at most $12g=p-$ 1 distinct eigenvalues).
By Dirichlet’s theorem on arithmetic progressions, we can find a prime $\ell$, not divisible by $\mathfrak{q}$, which is a primitive root modulo $p$. The reduction of $\tau$ modulo $\ell$ lands in $\operatorname{{\rm Gal}}(K(A[\ell])/K)$ and has order $p$. This allows us to apply [Goo21, Theorem 2.9] and deduce $E\coloneqq\operatorname{{\rm End}}^{0}(A)$ is a field.
As $p$ does not divide $[L:K]$, the element $\tau$ lies in the image of $G_{L}$. The decomposition $E\otimes\mathbb{Q}_{\ell}=\prod_{\lambda|\ell}E_{\lambda}$ induces a decomposition $V_{\ell}(A)=\prod_{\lambda|\ell}V_{\lambda}(A)$ giving representations $G_{L}\rightarrow\operatorname{{\rm GL}}_{n}(E_{\lambda})$ where $n=\frac{2g}{[E:\mathbb{Q}]}$, see §2 of [Rib76] for further details. Let $\tau_{\lambda}\in\operatorname{{\rm GL}}_{n}(E_{\lambda})$ be the projection of $\tau$ onto $V_{\lambda}(A)$. By the above, the eigenvalues of $\tau_{\lambda}$ are distinct primitive $p$-th roots of unity. Taking the trace of $\tau_{\lambda}$ we deduce $E_{\lambda}$ contains a subfield of $\mathbb{Q}(\zeta_{p})$ of degree $[E:\mathbb{Q}]$.
As $\mathbb{Q}(\zeta_{p})$ has a unique subfield $F$ of degree $[E:\mathbb{Q}]$ and for all, but possibly finitely many $\lambda$, $E_{\lambda}$ contains $F$, the Chebotarev Density Theorem implies $E$ contains $F$. But $[E:\mathbb{Q}]=[F:\mathbb{Q}]$, thus in other words, $E\subseteq\mathbb{Q}(\zeta_{p})$. ∎
By imposing a condition on the ray class groups of primes above $2$ in $K$, we deduce an explicit version of the above:
Theorem 2.4.
Let $A/K$ be an abelian variety of dimension $g\geq 1$ with $p=2g+1$ prime. Suppose $\operatorname{{\rm Gal}}(K(A[2])/K)\cong C_{p}$ and $p$ divides neither the class number of $K$, nor the multiplicative order of the residue field of any prime above 2. Then either
  • $p$ does not divide $[L:K]$ and $\operatorname{{\rm End}}^{0}(A)$ is a subfield of $\mathbb{Q}(\zeta_{p})$; or
  • $p$ divides $[L:K]$, $g\geq 3$ and $A$ is isogenous over $\bar{K}$ to the power of an absolutely simple abelian variety with complex multiplication by a proper subfield of $\mathbb{Q}(\zeta_{p})$.
In particular there are only finitely many possibilities for $\operatorname{{\rm End}}^{0}(A)$.
Proof.
Let $F/K$ be an odd degree abelian extension unramified outside of 2. Then by class field theory, any prime dividing $[F\colon K]$ divides either the class number of $K$, or the multiplicative order of the residue field of a prime above 2 in $K$.
It follows that $K(A[2])/K$ is ramified at some prime $\mathfrak{q}\nmid 2$. This allows us to apply the above proposition and conclude. ∎
The following lemma is well-known.
Lemma 2.5.
Let $A/K$ be an absolutely simple abelian variety with CM by a Galois extension $E/\mathbb{Q}$. Then $L=E^{*}K$ and $E\supseteq E^{*}$, where $E^{*}$ is the reflex field of $E$. Furthermore if $E/\mathbb{Q}$ is abelian, then $E=E^{*}$.
Proof.
As $E/\mathbb{Q}$ is Galois, the reflex field $E^{*}$ is a subfield of $E$ [Shi98, Prop. 28, pg 62]. Moreover, since $A$ is absolutely simple its CM type is primitive. Proposition 30 on page 65 of [Shi98] applies to tell us the endomorphism field $L$ equals $E^{*}K$. Finally, if $E/\mathbb{Q}$ is an abelian extension, then $E=E^{*}$ by Example (1) on page 63 of [Shi98]. ∎
Corollary 2.6.
Let $A/\mathbb{Q}$ be an abelian variety of dimension $g\geq 1$ with $p=2g+1$ prime. Suppose $\operatorname{{\rm Gal}}(\mathbb{Q}(A[2])/\mathbb{Q})\cong C_{p}$. Then either
  • $\operatorname{{\rm End}}^{0}(A)$ is a proper subfield of $\mathbb{Q}(\zeta_{p})$; or
  • $p\in\{7,11,19,43,67,163\}$ and $A$ is isogenous over $\bar{\mathbb{Q}}$ to the power of an elliptic curve with complex multiplication by $\mathbb{Q}(\sqrt{-p})$.
In particular there are only finitely many possibilities for $\operatorname{{\rm End}}^{0}(A)$.
Proof.
Clearly, we may apply Theorem 2.4. Suppose first $p$ divides $[L:\mathbb{Q}]$. Then $g\geq 3$ and $A$ is isogenous over $\bar{\mathbb{Q}}$ to the power of an absolutely simple abelian variety with complex multiplication by a proper (non-trivial) subfield $F$ of $\mathbb{Q}(\zeta_{p})$. Thus there is some minimal finite extension $K/\mathbb{Q}$ such that $A_{K}$ contains a simple abelian subvariety $B$ with CM by $F$. The above lemma applies to $B$ and shows $L$, the endomorphism field of $A$, contains $KF$.
Proposition 3.9 of [Goo21] (recalled at the beginning of this section) shows $\operatorname{{\rm Gal}}(L/\mathbb{Q})$ is an extension of $C_{p}$ by $C_{2}^{m}$ for some $m$. In particular, any non-trivial element of even order in $\operatorname{{\rm Gal}}(L/\mathbb{Q})$ has either order 2 or $2p$.
Since $p$ does not divide $[F:\mathbb{Q}]$, the cyclic quotient $\operatorname{{\rm Gal}}(F/\mathbb{Q})$ of $\operatorname{{\rm Gal}}(L/\mathbb{Q})$ has order two. Using $F$ is a CM field only ramified at $p$, we find $F=\mathbb{Q}(\sqrt{-p})$ where $p\equiv 3\mod{4}$. In particular $g$ is odd.
We show $F$ has class number one, which by the Baker–Heegner–Stark Theorem will conclude the proof in this case. As $F/\mathbb{Q}$ is a quadratic extension, $B$ is an elliptic curve. Thus by CM theory [Sil94, II §4 Thm. 4.1, pg 121] $KF$ contains $H$ the Hilbert class field of $F$. In particular, $L$ contains $H$, so $[H:F]$ divides $2^{m}p$ for some $m$. On the other hand, [FG18, Corollary 2.16] tells us every element of $\operatorname{{\rm Gal}}(H/F)$ has order dividing $g$. As $g$ is odd and less than $p$, we find $\operatorname{{\rm Gal}}(H/F)=1$ as claimed.
We now suppose $p$ does not divide $[L:\mathbb{Q}]$ and show $E\neq\mathbb{Q}(\zeta_{p})$. Suppose we had equality, then applying Lemma 2.5 we find $L=\mathbb{Q}(\zeta_{p})$. If $p=3$, then we apply Theorem 2.9 of [Goo21] and Theorem 2.1 to obtain a contradiction. Else, $p\geq 5$ and Proposition 3.9 of [Goo21] provides us with a contradiction, since $\operatorname{{\rm Gal}}(\mathbb{Q}(\zeta_{p})/\mathbb{Q})$ is not an elementary abelian $2$-group. ∎
Let us record some easy corollaries of Theorem 2.9 of [Goo21], Theorems 2.1, 2.6 and Lemma 2.5.
Corollary 2.7.
Let $A\colon y^{2}=f(x)$ be an elliptic curve defined over a number field with a real embedding. If $\operatorname{{\rm Gal}}(f)\cong C_{3}$, then $A$ does not have complex multiplication.
Corollary 2.8.
Let $A/\mathbb{Q}$ be an abelian surface. Suppose $\operatorname{{\rm Gal}}(\mathbb{Q}(A[2])/\mathbb{Q})\cong C_{5}$. Then either $\operatorname{{\rm End}}(A)=\mathbb{Z}$ or $\operatorname{{\rm End}}_{\mathbb{Q}}^{0}(A)=\operatorname{{\rm End}}^{0}(A)=\mathbb{Q}(\sqrt{5})$.
Example 2.9.
Both cases are possible. Indeed, examples of the first case are easily found. For the second case, note the jacobian $J$ of the hyperelliptic curve $y^{2}=x(x^{5}-4x^{4}+2x^{3}+5x^{2}-2x-1)$ has $\operatorname{{\rm End}}_{\mathbb{Q}}(J)=\operatorname{{\rm End}}(J)\cong\mathbb{Z}\left[\frac{1+\sqrt{5}}{2}\right]$ by [Wil00, Proposition 1] and the Galois group of $x^{5}-4x^{4}+2x^{3}+5x^{2}-2x-1$ has order 5.
We present a variant of Theorem 2.4 for abelian varieties over imaginary quadratic fields.
Theorem 2.10.
Let $g\geq 2$ be an integer and suppose $p=2g+1$ is prime. Let $K$ be an imaginary quadratic field of class number coprime to $p$. Let $A/K$ be an abelian variety of dimension $g$. Suppose $\operatorname{{\rm Gal}}(K(A[2])/K)\cong C_{p}$. Then either
  • $\operatorname{{\rm End}}^{0}(A)$ is a proper subfield of $\mathbb{Q}(\zeta_{p})$; or
  • $p\in\{7,11,19,43,67,163\}$ and $A$ is isogenous over $\bar{K}$ to the power of an elliptic curve with complex multiplication by $\mathbb{Q}(\sqrt{-p})$.
In particular there are only finitely many possibilities for $\operatorname{{\rm End}}^{0}(A)$.
Proof.
Clearly, we may apply Theorem 2.4. Suppose first $p$ divides $[L:K]$. Then $A$ is isogenous over $\bar{K}$ to the power of an absolutely simple abelian variety $B$ with complex multiplication by a proper subfield $F$ of $\mathbb{Q}(\zeta_{p})$. Lemma 2.5 applies to $B$ and shows $L$, the endomorphism field of $A$, contains $F$.
Proposition 3.9 of [Goo21] (recalled at the beginning of this section) shows $\operatorname{{\rm Gal}}(L/K)$ is an extension of $C_{p}$ by $C_{2}^{m}$ for some $m$. In particular, any non-trivial element of even order in $\operatorname{{\rm Gal}}(L/K)$ has either order 2 or $2p$. Since $p$ does not divide $[F:\mathbb{Q}]$, the cyclic quotient $\operatorname{{\rm Gal}}(FK/K)$ of $\operatorname{{\rm Gal}}(L/K)$ has order dividing two.
In turn we deduce $[F:\mathbb{Q}]$ divides $4$ and is equal to 4 only if $F\supsetneq K$. Being a CM field, $[F:\mathbb{Q}]=4$ would imply $F$ contains a real quadratic field. Whence $\operatorname{{\rm Gal}}(F/\mathbb{Q})\cong C_{2}\times C_{2}$ contradicting the fact $F\subseteq\mathbb{Q}(\zeta_{p})$. Thus $[F:\mathbb{Q}]=2$. As $F$ is a CM field only ramified at $p$, we deduce $F=\mathbb{Q}(\sqrt{-p})$ where $p\equiv 3\mod{4}$.
[…]
Thus either $H=\mathbb{Q}(\sqrt{-p})$ or $H=K(\sqrt{-p})$. Suppose the latter holds. Then, as $H$ is an unramified extension of $\mathbb{Q}(\sqrt{-p})$, which itself is only ramified at $p$, we see that $p$ is the only finite prime ramified in $K$. As $p$ is odd, $K$ is tamely ramified at $p$, so by the Kronecker-Weber Theorem $K\subseteq\mathbb{Q}(\zeta_{p})$. This implies $K=\mathbb{Q}(\sqrt{-p})$, since it is the unique quadratic field contained in $\mathbb{Q}(\zeta_{p})$. Thus $H=K(\sqrt{-p})=\mathbb{Q}(\sqrt{-p})$.
In particular, $\mathbb{Q}(\sqrt{-p})$ has class number one. As $p\equiv 3\mod{4}$ and $g\geq 2$, it follows, from the classification of imaginary quadratic fields of class number one, that $p\in\{7,11,19,43,67,163\}$.
We now suppose $p$ does not divide $[L:K]$ and show $E\neq\mathbb{Q}(\zeta_{p})$. Suppose we had equality, then by Lemma 2.5, $L=K(\zeta_{p})$. Proposition 3.9 of [Goo21] then implies $\operatorname{{\rm Gal}}(L/K)=\operatorname{{\rm Gal}}(K(\zeta_{p})/K)\cong\operatorname{{\rm Gal}}(\mathbb{Q}(\zeta_{p})/\mathbb{Q}(\zeta_{p})\cap K)$ has order at most $2$. Hence $[\mathbb{Q}(\zeta_{p}):\mathbb{Q}]$ divides $4$, being equal to 4 only if $K$ is contained in $\mathbb{Q}(\zeta_{p})$. But $[\mathbb{Q}(\zeta_{p}):\mathbb{Q}]=4$ implies $p=5$ and $\mathbb{Q}(\zeta_{5})$ does not have an imaginary subfield. Whence $[\mathbb{Q}(\zeta_{p}):\mathbb{Q}]=2$ and $p=3$. This in turn implies $g=1$ which we have ruled out by assumption. ∎
Remark 2.11.
The condition on the class number of $K$ cannot be removed. Indeed, the polynomial $f(x)=x^{5}-19x^{4}+107x^{3}+95x^{2}+88x-16$ has Galois group $D_{5}$ and its splitting field is the Hilbert class field of $K\coloneqq\mathbb{Q}(\sqrt{-131})$. The jacobian $J_{f}$ of the hyperelliptic curve defined by $y^{2}=f(x)$, has endomorphism algebra isomorphic to $\mathbb{Q}(\sqrt{13})$ and the group $\operatorname{{\rm Gal}}(K(J_{f}[2])/K)$ is isomorphic to $C_{5}$.
The above example raises the following question:
Question 2.12.
Let $K$ be an imaginary quadratic field of class number divisible by 5. Let $H$ be a degree 5 extension of $K$ contained in the Hilbert class field of $K$. What are the possible endomorphism algebras for abelian surfaces $A/\mathbb{Q}$ with $\mathbb{Q}(A[2])=H$?
Remark 2.13.
By combining Theorem 2.1, Theorem 2.10 of [Goo21] and [Shi98, Prop. 30, pg.65; Example (2), pg.64], we see that either $\operatorname{{\rm End}}^{0}(A)\cong\mathbb{Q}$ or $\mathbb{Q}(\sqrt{d})$ with $d\equiv 5\mod{8}$.
The above question then becomes: given $H$, which $d$ occur?
We finish this section with the following extension of [Goo21, Theorem 3.5].
Theorem 2.14.
Let $f\in K[x]$ be a polynomial of odd degree $n$ with Galois group isomorphic to a Frobenius group $G$ of order $n(n-1)$. Let $J_{f}$ be the jacobian associated to the hyperelliptic curve defined by $y^{2}=f(x)$.
Suppose $E={\rm End}^{0}(J_{f})$ is a number field of dimension $s$ over $\mathbb{Q}$. Then $E/\mathbb{Q}$ is Galois with ${\rm Gal}(E/\mathbb{Q})$ isomorphic to a quotient of $H$ the Frobenius complement of $G$.
Furthermore, $L/K$ is an extension of degree $s$ contained in $K(f)$, and as abstract groups, ${\rm Gal}(L/K)\cong{\rm Gal}(E/\mathbb{Q})$. Moreover, if $s=n-1$, then $L=EK$.
Finally, if $\operatorname{{\rm End}}(J_{f})$ is $2$-maximal, then $E$ is unramified at 2.
Proof.
This result is [Goo21, Theorem 3.5] with the extra assertion that if $s=n-1$, then $L=EK$. If $s=n-1$, then $J_{f}$ is an absolutely simple abelian variety with CM by $E$. We may therefore apply Lemma 2.5 to find $L=E^{*}K$ and $E^{*}\subseteq E$. Hence $[E^{*}:\mathbb{Q}]\geq[L:K]=s=[E:\mathbb{Q}]$. It follows we have equality $E^{*}=E$. ∎
Example 2.15.
For $f(x)=x^{5}-2$, it is well known the endomorphism algebra of $J_{f}$ is $\mathbb{Q}(\zeta_{5})$ and $\mathbb{Q}(\zeta_{5})$ is the unique degree 4 extension contained in the splitting field $\mathbb{Q}(f)$.
A more interesting example is given by the genus 2 curve with LMFDB label 28561.a.371293.1 first found in [BSS+16] (though we use an odd degree model computed using Magma [BCP97]). Here $f(x)=52x^{5}+104x^{4}+104x^{3}+52x^{2}+12x+1$, $\operatorname{{\rm Gal}}(f)\cong F_{5}$ and $J_{f}$ has CM by the number field $K$ defined by $x^{4}-x^{3}+2x^{2}+4x+3$. The unique degree 4 extension contained in $\mathbb{Q}(f)$ is given by $K$.
This field is totally inert at 2 (as predicted by [Goo21, Theorem 2.9]) and unramified outside 13. We note its class number is one, but in line with the theorems presented above, $\mathbb{Q}(f)/K$ is a degree 5 extension ramified only at 2.
3. An analogue of Theorem 1.2 for QM surfaces
Let $B=\mathbb{Q}+i\mathbb{Q}+j\mathbb{Q}+k\mathbb{Q}$ be an indefinite quaternion algebra generated by $i,j$ satisfying $i^{2}=D/m$, $j^{2}=m$, $ij=-ji$ and $k=ij$, where $D$ is a positive squarefree integer and $m|D$. We note that as $B$ is indefinite, $0>-m=\mathrm{n}(j)$, the reduced norm of $j$. As we shall only use the reduced norm, reduced trace and reduced discriminant, we shall refer to these simply as the norm, trace and discriminant respectively.
An order of $B$ is said to be hereditary if it has squarefree (reduced) discriminant (for alternate, equivalent, definitions see [DR04, Voi21]). We shall take $\mathcal{O}$ to be a hereditary order in $B$ of discriminant $D$.
An abelian surface $A/K$ is said to have quaternion multiplication by $\mathcal{O}$ if there is an isomorphism $\iota\colon\mathcal{O}\xrightarrow{\sim}\operatorname{{\rm End}}(A)$. We shall assume there is an element $\mu\in\mathcal{O}$ satisfying $\mu^{2}+D=0$. This is always the case when $\mathcal{O}$ is a maximal order [Voi21, 43.6.6, page 818] and furthermore such an element $\mu$ induces a principal polarisation on $A$ [Voi21, page 821]. Fix such an $A$, $\mu$ and $\iota$.
Following [DR04, Definition 3.3], we say $\chi\in B$ is a twist of $(\mathcal{O},\mu)$ if it lies in both $\mathcal{O}$ and the normaliser $N_{B^{*}}(\mathcal{O})$, has trace zero, $\mathrm{n}(\chi)$ divides $D$, and $\mu\chi=-\chi\mu$. The existence of such an element can be verified by a finite computation, we refer the reader to the remark at the bottom of page 9 of [DR04] for further details.
Owing in part to the fact $(\mathcal{O},\mu)$ has been fixed, there are only two possible values [^3] This follows from [Rot04, Lemmas 3.5 and 3.7]. Indeed, in the notation of the paper, we have $F=\mathbb{Q}$ and $D\neq 3$, as $D$ is divisible by an even number of primes. This forces $\omega_{odd}=1$. The cited lemmas in turn show $C_{2}\cong U_{0}\leq V_{0}\cong C_{2}\times C_{2}$. The non-trivial element of $U_{0}$ can be represented by $\mu=\sqrt{-D}$. This allows us to write $V_{0}=\langle[\mu],[\chi]\rangle$ where the representative $\chi$ may be taken to have reduced norm $m|D$. Moreover, any representative of $[\chi]$ has reduced norm $m$ up to a rational square. Likewise, representatives of the class $[\mu\chi]$ have reduced norm $D/m$ up to rational squares. $\mathrm{n}(\chi)$ can take and the product of these values is equal to $D$. We call $(\mathcal{O},\mu,\chi)$ a twisted principally polarised order, and say $(\mathcal{O},\mu,\chi)$ is of discriminant $D$ and norm $-\mathrm{n}(\chi)$.
As before, we let $L$ denote the endomorphism field of $A$. Dieulefait and Rotger [DR04, Theorem 3.4] showed $\operatorname{{\rm Gal}}(L/K)$ is isomorphic to one of the trivial group, $C_{2}$, or $C_{2}\times C_{2}$. Moreover, they proved in each case $\operatorname{{\rm End}}_{K}^{0}(A)$ is respectively isomorphic to $B$; one of $\mathbb{Q}(\mu),\mathbb{Q}(\chi),\mathbb{Q}(\mu\chi)$; or $\mathbb{Q}$.
This determines (and heavily restricts) the possible images of
$\operatorname{{\rm Gal}}(L/K)\rightarrow\operatorname{{\rm Aut}}(\operatorname{{\rm End}}(A))\cong N_{B^{*}}(\mathcal{O})/\mathbb{Q}^{*}$
arising from the natural action of $\operatorname{{\rm Gal}}(L/K)$ on $\operatorname{{\rm End}}(A)$.
To gain information on the intersection $L\cap K(A[2])$, we will use a description of orders of discriminant $D$ in $B$.
Lemma 3.1.
Suppose $2|D$ and $m\equiv 3\mod{4}$. Then
$\mathcal{O}=\mathbb{Z}+\frac{1}{2}(1+j+k)\mathbb{Z}+\frac{1}{2}(1+j-k)\mathbb{Z}+\frac{1}{2}(i+k)\mathbb{Z}$
is an order of discriminant $D$ in $B$. Moreover, any order $\mathcal{O}^{\prime}$ of discriminant $D$ which contains $\mathbb{Z}[1,mi,j,k]$, satisfies $\mathcal{O}^{\prime}\otimes\mathbb{Z}_{2}=\mathcal{O}\otimes\mathbb{Z}_{2}$.
Proof.
It is a routine calculation to show $\mathcal{O}$ is an order of discriminant $D$ (for an example, see [Vig80, pg. 85 - 86]).
As $m$ is odd we have $\mathbb{Z}[1,mi,j,k]\otimes\mathbb{Z}_{2}=\mathbb{Z}[1,i,j,k]\otimes\mathbb{Z}_{2}$. Thus we may assume $\mathcal{O}^{\prime}$ contains $\mathbb{Z}[1,i,j,k]$. The discriminant of $\mathbb{Z}[1,i,j,k]$ equals $4D$ and thus any element contained in $\mathcal{O}^{\prime}$ but not in $\mathbb{Z}[1,i,j,k]$ is of the form $\frac{1}{4}\alpha$ with $\alpha\in\mathbb{Z}[1,i,j,k]$.
Let us write such an element as $\frac{1}{4}\alpha=w+xi+yj+zk$. Considering the trace of $\frac{1}{4}\alpha$, which is an integer [Voi21, Cor. 10.3.3], we find $w\equiv 0\mod{2}$. Likewise as $\mathrm{n}(\frac{1}{4}\alpha)\in\mathbb{Z}$ [Voi21, Cor. 10.3.3], we have $\mathrm{n}(\alpha)\equiv 0\mod{16}$. Considering $\mathrm{n}(\alpha)\equiv 0\mod{2}$, we find $y\equiv 0\mod{2}$, which in turn combined with $\mathrm{n}(\alpha)\equiv 0\mod{4}$ implies $x^{2}+z^{2}\equiv 0\mod{4}$ and thus $x,z\equiv 0\mod{2}$.
Hence it suffices to check elements of the form $\frac{1}{2}(a+bi+cj+dk)$ with $a,b,c,d\in\{0,1\}$. We have $\mathrm{n}(a+bi+cj+dk)\equiv a-2b+c+2d\equiv 0\mod{4}$, from which we deduce $\frac{1}{2}(1+j+k)$, $\frac{1}{2}(1+i+j)$, $\frac{1}{2}(i+k)$ are the only integral such. In order for the discriminant of $\mathcal{O}^{\prime}$ to equal $D$ we see that all of these elements must belong to $\mathcal{O}^{\prime}$. ∎
Lemma 3.2.
Suppose $m\equiv 1\mod{4}$. Then
$\mathcal{O}=\mathbb{Z}+\frac{1}{2}(1+j)\mathbb{Z}+k\mathbb{Z}+\frac{1}{2}(i+k)\mathbb{Z}$
is an order of discriminant $D$ in $B$. Furthermore, if $D\equiv 1\mod{4}$ then
$\mathcal{O}_{1}=\mathbb{Z}+\frac{1}{2}(1+i)\mathbb{Z}+j\mathbb{Z}+\frac{1}{2}(j+k)\mathbb{Z}$
is an order of discriminant $D$ in $B$. Likewise, if $D\equiv 3\mod{4}$ then
$\mathcal{O}_{3}=\mathbb{Z}+\frac{1}{2}(1+k)\mathbb{Z}+j\mathbb{Z}+\frac{1}{2}(i+j)\mathbb{Z}$
is an order of discriminant $D$ in $B$.
Moreover, any order $\mathcal{O}^{\prime}$ of even discriminant $D$ which contains $\mathbb{Z}[1,mi,j,k]$ satisfies $\mathcal{O}^{\prime}\otimes\mathbb{Z}_{2}=\mathcal{O}\otimes\mathbb{Z}_{2}$. Any order $\mathcal{O}^{\prime}$ of discriminant $D\equiv t\mod{4}$ with $t\in\{1,3\}$, which contains $\mathbb{Z}[1,mi,j,k]$, satisfies either $\mathcal{O}^{\prime}\otimes\mathbb{Z}_{2}=\mathcal{O}\otimes\mathbb{Z}_{2}$ or $\mathcal{O}^{\prime}\otimes\mathbb{Z}_{2}=\mathcal{O}_{t}\otimes\mathbb{Z}_{2}$.
Proof.
The proof follows the same strategy as for Lemma 3.1. ∎
Theorem 3.3.
Let $A/K$ be an abelian surface with QM by a twisted principally polarised order $(\mathcal{O},\mu,\chi)$ of discriminant $D$ and norm $m$, where $\mathcal{O}\subseteq B$.
If $2|D$ and $m\equiv 3\mod{4}$, then $L\subseteq K(A[2])$.
Proof.
As $\operatorname{{\rm End}}(A)$ is a hereditary order, the results of [DR04, Theorem 3.4] apply to show a non-trivial element of $\operatorname{{\rm Gal}}(L/K)$ acts on $\operatorname{{\rm End}}(A)$ (possibly after scaling) by conjugation as one of $\mu,\chi$ or $\mu\chi$.
We look to determine the action of these elements on $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}/2\mathbb{Z}$. To do so we reduce a $\mathbb{Z}$-basis of $\operatorname{{\rm End}}(A)$ modulo $2$, thus we may work with $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}_{2}$ in place of $\operatorname{{\rm End}}(A)$. By considering the algebraic relations they satisfy, we may assume $\mu=k$, $\chi=j$ and $\mu\chi=mi$. Hence Lemma 3.1 allows us to take $\operatorname{{\rm End}}(A)=\mathbb{Z}+\frac{1}{2}(1+j+k)\mathbb{Z}+\frac{1}{2}(1+j-k)\mathbb{Z}+\frac{1}{2}(i+k)\mathbb{Z}$.
Let $X=\frac{1}{2}(1+j+k)$, $Y=\frac{1}{2}(1+j-k)$ and $Z=\frac{1}{2}(i+k)$. Let us examine the action of $i,j$ and $k$ on the basis of $\operatorname{{\rm End}}(A)$ given by $1,X,Y,Z$. Each of $i,j$ and $k$ fix $1$. For $i$ we have $iXi^{-1}=1-X$, $iYi^{-1}=1-Y$ and $iZi^{-1}=Z+Y-X$. For $j$ we have $jXj^{-1}=Y$, $jYj^{-1}=X$, $jZj^{-1}=-Z$. Looking at the coefficients, we see the action remains faithful on $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}/2\mathbb{Z}$. By proceeding as in the proof of Theorem 2.1 we are now done. ∎
Example 3.4.
The hereditary assumption is necessary. The following example shows not only $L$ need not be contained in $\mathbb{Q}(A[2])$ for a non-hereditary order, but also the result of Dieulefait and Rotger fails.
Let $J$ be the jacobian defined by the hyperelliptic curve associated to $y^{2}+y=6x^{5}+9x^{4}-x^{3}-3x^{2}$ with LMFDB label 20736.l.373248.1. This surface has QM by an order of (reduced) discriminant $6^{2}$ in $\left(\frac{2,3}{\mathbb{Q}}\right)$. In particular, the order is not hereditary. The endomorphism field has defining polynomial $x^{8}+4x^{6}+10x^{4}+24x^{2}+36$ and the two torsion field of $J$ is $\mathbb{Q}(\sqrt{2},\sqrt{3})$.
This curve, made easily available on the LMFDB [LMF], was first found in [BSS+16]. Computations linked to its endomorphism algebra were carried out in Magma using code from [CMSV19].
For ease of notation, let $F=K(A[2])$.
Theorem 3.5.
Let $A/K$ be an abelian surface with QM by a twisted principally polarised order $(\mathcal{O},\mu,\chi)$ of discriminant $D$ and norm $m$, where $\mathcal{O}\subseteq B$.
If $m\equiv 1\mod{4}$, then $\operatorname{{\rm End}}^{0}_{F}(A)$ contains at least one of $\mathbb{Q}(\sqrt{m}),\mathbb{Q}(\sqrt{-D})$ and $\mathbb{Q}(\sqrt{D/m})$. Moreover, if $D$ is even, then $\operatorname{{\rm End}}^{0}_{F}(A)$ contains at least one of $\mathbb{Q}(\sqrt{-D})$ and $\mathbb{Q}(\sqrt{D/m})$.
Proof.
As $\operatorname{{\rm End}}(A)$ is a hereditary order, the results of [DR04, Theorem 3.4] apply to show a non-trivial element of $\operatorname{{\rm Gal}}(L/K)$ acts on $\operatorname{{\rm End}}(A)$ (possibly after scaling) by conjugation as one of $\mu,\chi$ or $\mu\chi$.
We look to determine the action of these elements on $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}/2\mathbb{Z}$. To do so we reduce a $\mathbb{Z}$-basis of $\operatorname{{\rm End}}(A)$ modulo $2$, thus we may work with $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}_{2}$ in place of $\operatorname{{\rm End}}(A)$. By considering the algebraic relations they satisfy, we may assume $\mu=k$, $\chi=j$ and $\mu\chi=mi$. Lemma 3.2 allows us to take $\operatorname{{\rm End}}(A)$ equal to one of the three given orders. As these orders differ by permuting $i,j$ and $k$, we only give details for the case $\operatorname{{\rm End}}(A)=\mathbb{Z}+\frac{1}{2}(1+j)\mathbb{Z}+k\mathbb{Z}+\frac{1}{2}(i+k)\mathbb{Z}$.
Let $X=\frac{1}{2}(1+j)$, $Y=k$ and $Z=\frac{1}{2}(i+k)$. Let us examine the action of $i,j$ and $k$ on the basis of $\operatorname{{\rm End}}(A)$ given by $1,X,Y,Z$. Each of $i,j$ and $k$ fix $1$. For $i$ we have $iXi^{-1}=1-X$, $iYi^{-1}=-Y$ and $iZi^{-1}=Z-Y$. For $j$ we have $jXj^{-1}=X$, $jYj^{-1}=-Y$, $jZj^{-1}=-Z$. Looking at the coefficients, we see $j$ acts trivially on $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}/2\mathbb{Z}$ whereas $i$ and $k$ act by the same involution. Proceeding as in the proof of Theorem 2.1, we find at least one of $\mathbb{Q}(i)\cong\mathbb{Q}(\sqrt{D/m})$ and $\mathbb{Q}(k)\cong\mathbb{Q}(\sqrt{-D})$ is contained in $\operatorname{{\rm End}}_{F}^{0}(A)$. ∎
Example 3.6.
The jacobian $J$ of the hyperelliptic curve $y^{2}=-2x^{6}-12x^{5}-21x^{4}-10x^{3}-3x^{2}+6x+1$ has QM by the maximal order of $B=\left(\frac{3,5}{\mathbb{Q}}\right)$ which has discriminant 15. The endomorphism field of $J$ is $L=\mathbb{Q}(\sqrt{3},\sqrt{-3})$, we have $L\cap\mathbb{Q}(A[2])=\mathbb{Q}(\sqrt{-3})$ and $\operatorname{{\rm End}}^{0}_{\mathbb{Q}(\sqrt{-3})}(J)=\mathbb{Q}(\sqrt{-15})$. These calculations were performed in Magma using code from [CMSV19]. The curve comes from [LY20].

### v4

@@
 In this article we study the endomorphism algebras of abelian varieties $A$ defined over a given number field $K$ with large cyclic 2-torsion fields.
-A key step in doing so is to provide criteria for all the endomorphisms of $A$ to be defined over $K(A[2])$, the field generated by its 2-torsion.
-When $K=\mathbb{Q}$ and $\operatorname{{\rm Gal}}(\mathbb{Q}(A[2])/\mathbb{Q})$ is cyclic of prime order $p=2\dim(A)+1$, we prove the somewhat surprising result that there are only finitely many possibilities for the geometric endomorphism algebra $\operatorname{{\rm End}}(A)\otimes\mathbb{Q}$.
+A key step in doing so is to provide criteria for all the endomorphisms of $A$ to be defined over $K(A[2])$, the field extension generated by its 2-torsion.
+When $K=\mathbb{Q}$ and $\operatorname{{\rm Gal}}(\mathbb{Q}(A[2])/\mathbb{Q})$ is cyclic of prime order $p=2\dim(A)+1$, we prove that there are only finitely many possibilities for the geometric endomorphism algebra $\operatorname{{\rm End}}(A)\otimes\mathbb{Q}$.
 In fact, when $\dim(A)\not\in\{3,5,9,21,33,81\}$, we show $\operatorname{{\rm End}}(A)\otimes\mathbb{Q}$ is a proper subfield of the $p$-th cyclotomic field.
@@
 1. Introduction
-Faltings’ Isogeny Theorem tells us that for an abelian variety $A$, defined over a number field $K$, we can recover its ring of endomorphisms defined over $K$, $\operatorname{{\rm End}}_{K}(A)$, tensored by $\mathbb{Z}_{\ell}$ (where $\ell$ is a prime) from looking at the action of the absolute Galois group $G_{K}\coloneqq\operatorname{{\rm Gal}}(\bar{K}/K)$ on its $\ell$-adic Tate module $T_{\ell}(A)$.
+Faltings’ Isogeny Theorem [Fal83, p.
+360, Satz 4] tells us that for an abelian variety $A$, defined over a number field $K$, we can recover its ring of endomorphisms defined over $K$, $\operatorname{{\rm End}}_{K}(A)$, tensored by $\mathbb{Z}_{\ell}$ (where $\ell$ is a prime) from looking at the action of the absolute Galois group $G_{K}\coloneqq\operatorname{{\rm Gal}}(\overline{K}/K)$ on its $\ell$-adic Tate module $T_{\ell}(A)$.
 More precisely, Faltings’ Isogeny Theorem tells us that the natural injection
@@
 is not an isomorphism.
-However, when the action of $G_{K}$ on $A[\ell]$ is particularly ‘large’ one might hope that $\operatorname{{\rm End}}_{K}(A[\ell])\;(=\operatorname{{\rm End}}_{\bar{K}}(A[\ell])^{G_{K}})$ is small enough to place significant restrictions on $\operatorname{{\rm End}}_{K}(A)\otimes\mathbb{Z}/\ell\mathbb{Z}$ and hence $\operatorname{{\rm End}}_{K}(A)$.
+However, when the action of $G_{K}$ on $A[\ell]$ is particularly ‘large’ one might hope that $\operatorname{{\rm End}}_{K}(A[\ell])\;(=\operatorname{{\rm End}}_{\overline{K}}(A[\ell])^{G_{K}})$ is small enough to place significant restrictions on $\operatorname{{\rm End}}_{K}(A)\otimes\mathbb{Z}/\ell\mathbb{Z}$ and hence $\operatorname{{\rm End}}_{K}(A)$.
 Zarhin has had much success with carrying this out for jacobians of hyperelliptic curves (with $\ell=2$), see for example [Zar00, Zar01, Zar05].
@@
 In the cases considered by Zarhin, all of the Galois groups are insoluble and act 2-transitively on the roots of $f$.
-In [Goo21] the author considered restrictions on $\operatorname{{\rm End}}_{K}(A)$ and $\operatorname{{\rm End}}(A)\coloneqq\operatorname{{\rm End}}_{\bar{K}}(A)$ when $\operatorname{{\rm Gal}}(K(A[\ell])/K)$ contained merely an element of large prime order, with no restrictions on $K$.
-However, considering the case of elliptic curves, for example, where the possible endomorphism algebras depend solely on $[K:\mathbb{Q}]$, it is natural to impose conditions on $K$.
+In [Goo21] the author considered restrictions on $\operatorname{{\rm End}}_{K}(A)$ and $\operatorname{{\rm End}}(A)\coloneqq\operatorname{{\rm End}}_{\overline{K}}(A)$ when $\operatorname{{\rm Gal}}(K(A[\ell])/K)$ contained merely an element of large prime order, with no restrictions on $K$.
+However, considering the case of elliptic curves, for example, where the possible endomorphism algebras depend solely on $K\colon\mathbb{Q}$], it is natural to impose conditions on $K$.
 The main advantage in doing so is that this allows for more arithmetic input, whereas the proofs of the above results are heavily representation theoretic.
@@
 In this way, this paper may be seen as a natural continuation of [Goo21].
-As an example, we obtain the somewhat surprising [^1] At least in the author’s opinion. result (where $\zeta_{n}$ denotes a primitive $n$-th root of unity):
+As an example, we obtain (where $\zeta_{n}$ denotes a primitive $n$-th root of unity):
 Theorem 1.1 ($\subset$ Corollary 2.6 + Theorem 2.10).
@@
   • $\operatorname{{\rm End}}(A)\otimes\mathbb{Q}$ is a proper subfield of $\mathbb{Q}(\zeta_{p})$; or
-  • $p\in\{7,11,19,43,67,163\}$ and $A$ is isogenous over $\bar{K}$ to the power of an elliptic curve with complex multiplication by $\mathbb{Q}(\sqrt{-p})$.
+  • $p\in\{7,11,19,43,67,163\}$ and $A$ is isogenous over $\overline{K}$ to the power of an elliptic curve with complex multiplication by $\mathbb{Q}(\sqrt{-p})$.
 In particular there are only finitely many possibilities for $\operatorname{{\rm End}}^{0}(A)$.
@@
 To make this rigorous, we need to study the minimal extension $L/K$ over which $A$ obtains all its endomorphisms.
-This field was christened in [GK17] as the endomorphism field of $A$
-An important result in understanding this field was obtained by Silverberg, who showed it is contained in $K(A[n])$ for $n\geq 3$ [Sil92, Theorem 2.4].
-Bounds on $[L\colon K]$ were then studied in [GK17] and also in [FKRS12, FKS21] via its connection to Sato-Tate groups.
+This field was christened in [GK17] as the endomorphism field of $A$ .
+An important result in understanding this field was obtained by Silverberg, who showed it is contained in $K(A[n])$ for $n\geq 3$ [Sil92, Thm.
+2.4].
+Bounds on $L\colon K]$ were then studied in [GK17] and also in [FKRS12, FKS25] via its connection to Sato-Tate groups.
 However, as eluded to above, we wish to be able to bound $L$ relative to $K(A[2])$. The following theorem allows us to do so (see also Theorems 3.3, 3.5 for an analogue in the setting of QM surfaces):
-Theorem 1.2 ($=$ Theorem 2.1).
-Suppose $E=\operatorname{{\rm End}}(A)\otimes\mathbb{Q}$ is a (finite) Galois extension of $\mathbb{Q}$ and $L\nsubseteq K(A[2])$. The following hold:
-  • $\operatorname{{\rm Gal}}(E/\mathbb{Q})$ has a non-trivial normal elementary abelian 2-subgroup;
-  • if $\operatorname{{\rm End}}(A)$ is 2-maximal in $E$, then $2$ is wildly ramified in $E/\mathbb{Q}$.
-In particular, if $E/\mathbb{Q}$ is Galois, $\operatorname{{\rm End}}(A)$ is a 2-maximal order and $2$ is not wildly ramified, then $L\subseteq K(A[2])$.
-Recall that an order $\mathcal{O}\subseteq\operatorname{{\rm End}}(A)\otimes\mathbb{Q}$ is said to be 2-maximal if for any other order $\mathcal{O}^{\prime}\subseteq\operatorname{{\rm End}}(A)\otimes\mathbb{Q}$ satisfying $\mathcal{O}\otimes\mathbb{Z}_{2}\subseteq\mathcal{O}^{\prime}\otimes\mathbb{Z}_{2}$ we have $\mathcal{O}\otimes\mathbb{Z}_{2}=\mathcal{O}^{\prime}\otimes\mathbb{Z}_{2}$.
+Theorem 1.2 (== Theorem 2.1).
+Suppose $E\cong\operatorname{{\rm End}}^{0}(A)$ is a (finite) Galois extension of $\mathbb{Q}$ and $L\nsubseteq K(A[2])$. Then $E\colon\mathbb{Q}$] is even.
+If moreover $\operatorname{{\rm End}}(A)$ is 2-maximal in $E$, then $\operatorname{{\rm Gal}}(E/\mathbb{Q})$ has a non-trivial normal elementary abelian 2-subgroup which is contained in the intersection of inertia groups $\cap_{\mathfrak{p}|2}I_{\mathfrak{p}}$ where $\mathfrak{p}$ runs over all primes above $2$ in $E$.
+In particular, if $E/\mathbb{Q}$ is Galois, $\operatorname{{\rm End}}(A)$ is a 2-maximal order and $2$ is not wildly ramified in $E$, then $L\subseteq K(A[2])$.
+Recall that an order $\mathcal{O}$ in a number field $E$ is said to be 2-maximal if any of the equivalent conditions hold: its index in the maximal order $\mathcal{O}_{E}$ is odd; $\mathcal{O}\otimes\mathbb{Z}_{2}=\mathcal{O}_{E}\otimes\mathbb{Z}_{2}$; or $\mathcal{O}\otimes\mathbb{Z}_{(2)}=\mathcal{O}_{E}\otimes\mathbb{Z}_{(2)}$ [Voi21, Lem.
+9.5.3, Lem.
+9.6.7].
 Example 2.2 shows that the condition $\operatorname{{\rm End}}(A)$ is 2-maximal cannot be removed.
-Finally, we would like to end the introduction with a challenge [^2] See also Question 2.12 for the reader which we hope will provide some impetus for studying moduli spaces of abelian surfaces.
-To do so let us state the following corollary of Theorems 1.1,1.2:
-Corollary 1.3 ($=$ Corollary 2.8).
+Finally, let us underline the following corollary of Theorems 1.1,1.2.
+Corollary 1.3 (== Corollary 2.8).
 Let $A/\mathbb{Q}$ be an abelian surface.
@@
 Both cases are possible, see Example 2.9.
-Challenge 1.4.
-Give an independent proof of Corollary 1.3 using moduli spaces of abelian surfaces.
 Acknowledgements.
-The author thanks Tim Dokchitser, Francesc Fité, Xavier Guitart, Jef Laga, Jeremy Rickard, Víctor Rotger, Ciaran Schembri, John Voight and David Zywina for useful comments, questions and discussions.
+The author thanks Tim Dokchitser, Francesc Fité, Xavier Guitart, Jef Laga, Jeremy Rickard, Víctor Rotger, Ciaran Schembri, John Voight, David Zywina and the anonymous referee for useful comments, questions and discussions.
 The author thanks the University of Bristol and the Université Clermont Auvergne for their hospitality during the period this paper was written.
 Where the author was funded by the EPSRC (grant number EP/N509619/1) and an early career fellowship from the LMS respectively.
-Finally, the author thanks Universitat de Barcelona where final revisions to this paper were made.
+Finally, the author thanks Universitat de Barcelona where final revisions to this paper were made whilst he was supported by the Spanish Ministry of Science and Innovation via the grant “Abelian varieties, L-functions, and rational points” (code PID2022-137605NB-I00).
 Notation.
-Throughout $A$ will denote an abelian variety of dimension $g$ defined over a number field $K$ (often $\mathbb{Q}$).
-Moreover, we denote by $L$ the smallest extension of $K$ over which all the endomorphisms of $A$ are defined.
-We write $\operatorname{{\rm End}}_{F}(A)$ for the ring of endomorphisms of $A$ defined over $F/K$, $\operatorname{{\rm End}}(A)\coloneqq\operatorname{{\rm End}}_{\bar{K}}(A)$ and $\operatorname{{\rm End}}_{F}^{0}(A)\coloneqq\operatorname{{\rm End}}_{F}(A)\otimes\mathbb{Q}$.
+Throughout $A$ will denote an abelian variety of dimension $g$ defined over a number field $K$.
+Moreover, fixing an algebraic closure of $\overline{K}$ of $K$, we denote by $L$ the smallest extension of $K$ contained in $\overline{K}$ over which all the endomorphisms of $A$ are defined.
+For an extension $F/K$ contained in $\overline{K}$, we write $\operatorname{{\rm End}}_{F}(A)$ for the ring of endomorphisms of $A$ defined over $F$, $\operatorname{{\rm End}}(A)\coloneqq\operatorname{{\rm End}}_{\overline{K}}(A)$ and $\operatorname{{\rm End}}_{F}^{0}(A)\coloneqq\operatorname{{\rm End}}_{F}(A)\otimes\mathbb{Q}$.
+Letting $G_{F}$ denote the absolute Galois group of $F$, we have $\operatorname{{\rm End}}(A)^{G_{F}}=\operatorname{{\rm End}}_{F}(A)$.
+We write $\mathrm{Cl}(F)$ for the class group of $F$ and $\mathrm{Cl}(F)$ for its order.
 We denote by $C_{n},D_{n},F_{n}$ the cyclic group of order $n$, the dihedral group of order $2n$, and the Frobenius group isomorphic to $\mathbb{F}_{n}\rtimes\mathbb{F}_{n}^{*}$ (for $n$ a prime power) respectively.
@@
 2. Proof of the main theorems
-Silverberg has shown the endomorphism field $L/K$ of an abelian variety $A/K$ is contained in the $n$-torsion field for every $n\geq 3$ [Sil92, Theorem 2.4].
-The case $n=4$ implies $\operatorname{{\rm Gal}}(L/L\cap K(A[2]))$ is an elementary abelian 2-group, since $\operatorname{{\rm Gal}}(K(A[4])/K(A[2]))$ is too, for details see [Goo21, Proposition 3.9].
+Silverberg has shown the endomorphism field $L$ of an abelian variety $A/K$ is contained in the $n$-torsion field for every $n\geq 3$ [Sil92, Thm.
+2.4].
+The case $n=4$ implies $\operatorname{{\rm Gal}}(L/L\cap K(A[2]))$ is an elementary abelian 2-group, since $\operatorname{{\rm Gal}}(K(A[4])/K(A[2]))$ is too, for details see [Goo21, Prop.
+3.9].
 Below we provide conditions on $\operatorname{{\rm End}}(A)$ which guarantee $L\subseteq K(A[2])$.
 Theorem 2.1.
-Suppose $E=\operatorname{{\rm End}}^{0}(A)$ is a (finite) Galois extension of $\mathbb{Q}$ and $L\nsubseteq K(A[2])$. The following hold:
-  • $\operatorname{{\rm Gal}}(E/\mathbb{Q})$ has a non-trivial normal elementary abelian 2-subgroup;
-  • if $\operatorname{{\rm End}}(A)$ is 2-maximal in $E$, then $2$ is wildly ramified in $E/\mathbb{Q}$.
-In particular, if $E/\mathbb{Q}$ is Galois, $\operatorname{{\rm End}}(A)$ is a 2-maximal order and $2$ is not wildly ramified, then $L\subseteq K(A[2])$.
+Suppose $E\cong\operatorname{{\rm End}}^{0}(A)$ is a (finite) Galois extension of $\mathbb{Q}$ and $L\nsubseteq K(A[2])$. Then $E\colon\mathbb{Q}$] is even.
+If moreover $\operatorname{{\rm End}}(A)$ is 2-maximal in $E$, then $\operatorname{{\rm Gal}}(E/\mathbb{Q})$ has a non-trivial normal elementary abelian 2-subgroup which is contained in the intersection of inertia groups $\cap_{\mathfrak{p}|2}I_{\mathfrak{p}}$ where $\mathfrak{p}$ runs over all primes above $2$ in $E$.
+In particular, if $E/\mathbb{Q}$ is Galois, $\operatorname{{\rm End}}(A)$ is a 2-maximal order and $2$ is not wildly ramified in $E$, then $L\subseteq K(A[2])$.
 Proof.
-If $\operatorname{{\rm Gal}}(L/K)$ acts faithfully on $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}/2\mathbb{Z}$, then we may recuperate Lemma 2.1 in [Sil92] with $N=2$, $\Lambda=\operatorname{{\rm End}}(A)$ and $G$ the image of $\operatorname{{\rm Gal}}(L/K)$ in $\operatorname{{\rm Aut}}(\Lambda)$.
-Following through the proofs of Propositions 2.2, 2.3 and Theorem 2.4 in [Sil92] shows $L\subseteq K(A[2])$.
-Hence we may suppose the action of $\operatorname{{\rm Gal}}(L/K)$ on $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}/2\mathbb{Z}$ has a non-trivial element $\gamma$ in its kernel.
-Since $\operatorname{{\rm Gal}}(L/K)$ acts on $\operatorname{{\rm End}}(A)$ by field automorphisms of $\operatorname{{\rm End}}^{0}(A)$, we may view its image as a subgroup of $\operatorname{{\rm Gal}}(E/\mathbb{Q})$.
-Identify $\gamma$ with its image in $\operatorname{{\rm Gal}}(E/\mathbb{Q})$.
-Let $d$ be the rank of $\operatorname{{\rm End}}(A)$ as a free $\mathbb{Z}$-module and view $\operatorname{{\rm Gal}}(E/\mathbb{Q})$ as a (finite) subgroup of $\operatorname{{\rm GL}}_{d}(\mathbb{Z})$ via its action on $\operatorname{{\rm End}}(A)$.
-The reduction map $\operatorname{{\rm GL}}_{d}(\mathbb{Z})\rightarrow\operatorname{{\rm GL}}_{d}(\mathbb{Z}/4\mathbb{Z})$ is injective on finite groups and the kernel of $\operatorname{{\rm GL}}_{d}(\mathbb{Z}/4\mathbb{Z})\rightarrow\operatorname{{\rm GL}}_{d}(\mathbb{Z}/2\mathbb{Z})$ is an elementary abelian $2$-group (for more details see [Goo21, Prop.
-3.9]).
-Thus the kernel of the map $\operatorname{{\rm Gal}}(E/\mathbb{Q})\rightarrow\operatorname{{\rm Aut}}(\operatorname{{\rm End}}(A)\otimes\mathbb{Z}/2\mathbb{Z})\cong\operatorname{{\rm GL}}_{d}(\mathbb{Z}/2\mathbb{Z})$ is an elementary abelian $2$-group which is non-trivial since it contains $\gamma$.
+Suppose $\operatorname{{\rm Gal}}(L/K)$ acts faithfully on $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}/2\mathbb{Z}$.
+Recall that the natural map $\operatorname{{\rm End}}(A)\rightarrow\operatorname{{\rm End}}(A[2])$ induces an injection $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}/2\mathbb{Z}\hookrightarrow\operatorname{{\rm End}}(A[2])$.
+The group $\operatorname{{\rm Gal}}(\overline{K}/K(A[2]))$ acts trivially on $\operatorname{{\rm End}}(A[2])$ and hence trivially on $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}/2\mathbb{Z}$ as well.
+Since $\operatorname{{\rm Gal}}(L/K)$ acts faithfully on $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}/2\mathbb{Z}$, we find $L\subseteq K(A[2])$.
+As we are assuming $L\not\subseteq K(A[2])$, we deduce that the action of $\operatorname{{\rm Gal}}(L/K)$ on $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}/2\mathbb{Z}$ has a non-trivial element $\gamma$ in its kernel.
+The action of $\operatorname{{\rm Gal}}(L/K)$ on $\operatorname{{\rm End}}(A)$ is faithful, allowing us to view $\operatorname{{\rm Gal}}(L/K)$ as a subgroup of $\operatorname{{\rm Aut}}_{\mathbb{Z}}(\operatorname{{\rm End}}(A))\cong\operatorname{{\rm GL}}_{d}(\mathbb{Z})$ where $d$ is the rank of $\operatorname{{\rm End}}(A)$ viewed as a free $\mathbb{Z}$-module.
+The action also extends to field automorphisms of $\operatorname{{\rm End}}^{0}(A)\cong E$, thus we may also view $\operatorname{{\rm Gal}}(L/K)$ as a subgroup of $\operatorname{{\rm Gal}}(E/\mathbb{Q})$.
+Any finite subgroup in the kernel of the reduction map $\operatorname{{\rm GL}}_{d}(\mathbb{Z})\rightarrow\operatorname{{\rm GL}}_{d}(\mathbb{Z}/2\mathbb{Z})$ is an elementary abelian 2-group [GL06, p.
+151, Lem.
+9].
+As $\operatorname{{\rm Aut}}_{\mathbb{Z}/2\mathbb{Z}}(\operatorname{{\rm End}}(A)\otimes\mathbb{Z}/2\mathbb{Z})\cong\operatorname{{\rm GL}}_{d}(\mathbb{Z}/2\mathbb{Z})$, we deduce that $\{\sigma\in\operatorname{{\rm Gal}}(L/K)|(\sigma-1)\operatorname{{\rm End}}(A)\subseteq 2\cdot\operatorname{{\rm End}}(A)\}$ is an elementary abelian $2$-group which is non-trivial since it contains $\gamma$.
+Hence $E\colon\mathbb{Q}$] is divisible by 2.
 Assume now $\operatorname{{\rm End}}(A)$ is a 2-maximal order in $E$.
-As $\gamma$ acts trivially on the order $\operatorname{{\rm End}}(A)$ modulo $2$, it also acts trivially on $\operatorname{{\rm End}}(A)$ modulo every prime above $2$ in $E$.
-Since $\operatorname{{\rm End}}(A)$ is 2-maximal, we find $\gamma$ belongs to the inertia group $I_{\mathfrak{p}}$ for every prime $\mathfrak{p}$ above 2 in $E$.
-By the above $\gamma$ has order two, and hence $2$ is wildly ramified in $E/\mathbb{Q}$.
+Hence $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}_{2}\cong\mathcal{O}_{E}\otimes\mathbb{Z}_{2}$ where $\mathcal{O}_{E}$ is the maximal order of $E$.
+As $\operatorname{{\rm Gal}}(E/\mathbb{Q})$ acts on $\mathcal{O}_{E}$, it also acts on $\mathcal{O}_{E}\otimes\mathbb{Z}_{2}$ and thus also on $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}_{2}$ via the above isomorphism.
+The kernel $H$ of the composite map $\operatorname{{\rm Gal}}(E/\mathbb{Q})\hookrightarrow\operatorname{{\rm Aut}}_{\mathbb{Z}_{2}}(\operatorname{{\rm End}}(A)\otimes\mathbb{Z}_{2})\rightarrow\operatorname{{\rm Aut}}_{\mathbb{Z}/2\mathbb{Z}}(\operatorname{{\rm End}}(A)\otimes\mathbb{Z}/2\mathbb{Z})$ is an elementary abelian 2 group [GL06, p.
+151, Lem.
+9] [^1] The statement they give does not apply directly to $\mathbb{Z}_{2}$, however the proof carries over verbatim., which is non-trivial since it contains $\gamma$.
+Write $2\mathcal{O}_{E}=\prod_{i}\mathfrak{p}^{e_{i}}_{i}$ where each $\mathfrak{p}_{i}$ is a prime ideal.
+We have that every element of $H$ acts trivially on $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}/2\mathbb{Z}\cong\mathcal{O}_{E}\otimes\mathbb{Z}/2\mathbb{Z}\cong\mathcal{O}_{E}/2\mathcal{O}_{E}\cong\prod_{i}\mathcal{O}_{E}/\mathfrak{p}^{e_{i}}_{i}$.
+In particular, every element of $H$ acts trivially on each $\mathcal{O}/\mathfrak{p}^{e_{i}}_{i}$, so $H$ is contained in the inertia group $I_{\mathfrak{p}_{i}}$ for every prime $\mathfrak{p}_{i}$ above 2 in $E$.
 ∎
@@
 Let $A/K$ be an abelian variety of dimension $g\geq 1$ with $p=2g+1$ prime.
-Suppose $\mathfrak{q}$ is a prime of bad reduction for $A$ and there is an element of order $p$ in the image of the inertia group $I_{\mathfrak{q}}$ on $A[\ell]$ for some $\mathfrak{q}\nmid\ell$.
+Suppose $\mathfrak{q}$ is a prime of bad reduction for $A$ and there is an element of order $p$ in the image of the inertia group $I_{\mathfrak{q}}$ in $\operatorname{{\rm Gal}}(K(A[\ell^{\prime}])/K)$ for some prime $\ell^{\prime}\neq p$ and $\mathfrak{q}\nmid\ell^{\prime}$.
 Then either
-  • $p$ does not divide $[L:K]$ and $\operatorname{{\rm End}}^{0}(A)$ is a subfield of $\mathbb{Q}(\zeta_{p})$; or
-  • $p$ divides $[L:K]$, $g\geq 3$ and $A$ is isogenous over $\bar{K}$ to the power of an absolutely simple abelian variety with complex multiplication by a proper subfield of $\mathbb{Q}(\zeta_{p})$.
+  • $p$ does not divide $L\colon K]$ and $\operatorname{{\rm End}}^{0}(A)$ is isomorphic to a subfield of $\mathbb{Q}(\zeta_{p})$; or
+  • $p$ divides $L\colon K]$, $g\geq 3$ and $A$ is isogenous over $\overline{K}$ to the power of an absolutely simple abelian variety with complex multiplication by a proper subfield of $\mathbb{Q}(\zeta_{p})$.
 In particular there are only finitely many possibilities for $\operatorname{{\rm End}}^{0}(A)$.
 Proof.
-Suppose $p$ divides $[L:\mathbb{Q}]$.
-Then by Theorem 2.5 of [Goo21], $g\geq 3$ and $A$ is isogenous over $\bar{K}$ to the power of an absolutely simple abelian variety with complex multiplication by a proper subfield $F$ of $\mathbb{Q}(\zeta_{p})$.
-Hence we may suppose $p$ does not divide $[L:K]$.
-The criterion of Néron-Ogg-Shafarevich implies that the image of the inertia group $I_{\mathfrak{q}}$ on the Tate module $T_{\ell}(A)$ for any prime $\ell$, not divisible by $\mathfrak{q}$, contains an element $\tau$ of order $p$.
-Furthermore, as the trace of $\tau$ is an integer [ST68, Thm.
-2] its eigenvalues are the primitive $p$-th roots of unity (owing to the fact $\tau$ has at most $12g=p-$ 1 distinct eigenvalues).
-By Dirichlet’s theorem on arithmetic progressions, we can find a prime $\ell$, not divisible by $\mathfrak{q}$, which is a primitive root modulo $p$.
-The reduction of $\tau$ modulo $\ell$ lands in $\operatorname{{\rm Gal}}(K(A[\ell])/K)$ and has order $p$.
-This allows us to apply [Goo21, Theorem 2.9] and deduce $E\coloneqq\operatorname{{\rm End}}^{0}(A)$ is a field.
-As $p$ does not divide $[L:K]$, the element $\tau$ lies in the image of $G_{L}$.
-The decomposition $E\otimes\mathbb{Q}_{\ell}=\prod_{\lambda|\ell}E_{\lambda}$ induces a decomposition $V_{\ell}(A)=\prod_{\lambda|\ell}V_{\lambda}(A)$ giving representations $G_{L}\rightarrow\operatorname{{\rm GL}}_{n}(E_{\lambda})$ where $n=\frac{2g}{[E:\mathbb{Q}]}$, see §2 of [Rib76] for further details.
-Let $\tau_{\lambda}\in\operatorname{{\rm GL}}_{n}(E_{\lambda})$ be the projection of $\tau$ onto $V_{\lambda}(A)$.
+Suppose $p$ divides $L\colon\mathbb{Q}$].
+Then by [Goo21, Thm.
+2.5], $g\geq 3$ and $A$ is isogenous over $\overline{K}$ to the power of an absolutely simple abelian variety with complex multiplication by a proper subfield $F$ of $\mathbb{Q}(\zeta_{p})$.
+Hence we may suppose $p$ does not divide $L\colon K]$.
+Let $\ell$ be a prime and let $\rho_{\ell}\colon G_{K}\rightarrow\operatorname{{\rm Aut}}_{\mathbb{Z}_{\ell}}(T_{\ell}(A))$ denote the representation of $G_{K}$ arising from its action on the Tate module $T_{\ell}(A)$.
+By [GR72, Thm.
+4.3], for any $\sigma\in I_{\mathfrak{q}}$ and prime $\ell\nmid N(\mathfrak{q})$, the characteristic polynomial of $\rho_{\ell}(\sigma)$ has rational coefficients independent of $\ell$.
+By assumption, there exists a prime $\ell^{\prime}\neq p$ such that the image of $I_{\mathfrak{q}}$ in $\operatorname{{\rm Gal}}(K(A[\ell^{\prime}])/K)$ contains an element of order $p$.
+In particular, as $\ell^{\prime}\neq p$, there exists an element $\tau_{\ell^{\prime}}$ in the image of $I_{\mathfrak{q}}$ in $\operatorname{{\rm Aut}}_{\mathbb{Z}_{\ell^{\prime}}}(T_{\ell^{\prime}}(A))$ of order $p$.
+As the trace of $\tau_{\ell^{\prime}}$ is rational its eigenvalues are the primitive $p$-th roots of unity, owing to the fact $\tau_{\ell^{\prime}}$ has at most 2g=p12g=p-1 distinct eigenvalues.
+It follows that the characteristic polynomial of $\tau_{\ell^{\prime}}$ is the $p$-th cyclotomic polynomial.
+Let $\tau\in I_{\mathfrak{q}}$ be such that $\rho_{\ell^{\prime}}(\tau)=\tau_{\ell^{\prime}}$.
+We deduce that for any prime $\ell\nmid N(\mathfrak{q})$, the characteristic polynomial of $\tau_{\ell}\coloneqq\rho_{\ell}(\tau)$ is the $p$-th cyclotomic polynomial.
+By Dirichlet’s theorem on arithmetic progressions, we can find a prime $\ell^{\prime\prime}\nmid N(\mathfrak{q})$, which is a primitive root modulo $p$.
+The reduction of $\tau_{\ell^{\prime\prime}}$ modulo $\ell^{\prime\prime}$ lands in $\operatorname{{\rm Gal}}(K(A[\ell^{\prime\prime}])/K)$ and has order $p$ thanks to the form of its characteristic polynomial.
+This allows us to apply [Goo21, Thm.
+2.9] and deduce $E\coloneqq\operatorname{{\rm End}}^{0}(A)$ is a field.
+As $p$ does not divide $L\colon K]$, for any prime $\ell\nmid N(\mathfrak{q})$, the element $\tau_{\ell}$ lies in the image of $G_{L}$ in $\operatorname{{\rm Aut}}_{\mathbb{Z}_{\ell}}(T_{\ell}(A))$.
+The decomposition $E\otimes\mathbb{Q}_{\ell}=\prod_{\lambda|\ell}E_{\lambda}$ induces a decomposition $V_{\ell}(A)=\prod_{\lambda|\ell}V_{\lambda}(A)$ giving representations $G_{L}\rightarrow\operatorname{{\rm GL}}_{n}(E_{\lambda})$ where $n=\frac{2g}{[E\colon\mathbb{Q}]}$, see [Rib76, Ch.
+II] for further details.
+Let $\tau_{\lambda}\in\operatorname{{\rm GL}}_{n}(E_{\lambda})$ be the projection of $\tau_{\ell}$ onto $V_{\lambda}(A)$.
 By the above, the eigenvalues of $\tau_{\lambda}$ are distinct primitive $p$-th roots of unity.
-Taking the trace of $\tau_{\lambda}$ we deduce $E_{\lambda}$ contains a subfield of $\mathbb{Q}(\zeta_{p})$ of degree $[E:\mathbb{Q}]$.
-As $\mathbb{Q}(\zeta_{p})$ has a unique subfield $F$ of degree $[E:\mathbb{Q}]$ and for all, but possibly finitely many $\lambda$, $E_{\lambda}$ contains $F$, the Chebotarev Density Theorem implies $E$ contains $F$.
-But $[E:\mathbb{Q}]=[F:\mathbb{Q}]$, thus in other words, $E\subseteq\mathbb{Q}(\zeta_{p})$.
+Taking the trace of $\tau_{\lambda}$ we deduce $E_{\lambda}$ contains a subfield of $\mathbb{Q}(\zeta_{p})$ of degree $E\colon\mathbb{Q}$].
+As $\mathbb{Q}(\zeta_{p})$ has a unique subfield $F$ of degree $E\colon\mathbb{Q}$] and for all, but possibly finitely many primes $\lambda$, $E_{\lambda}$ contains $F$, the Chebotarev Density Theorem implies $E$ contains $F$.
+But $E\colon\mathbb{Q}]=[F\colon\mathbb{Q}$], so in other words, $E\subseteq\mathbb{Q}(\zeta_{p})$.
 ∎
@@
 Then either
-  • $p$ does not divide $[L:K]$ and $\operatorname{{\rm End}}^{0}(A)$ is a subfield of $\mathbb{Q}(\zeta_{p})$; or
-  • $p$ divides $[L:K]$, $g\geq 3$ and $A$ is isogenous over $\bar{K}$ to the power of an absolutely simple abelian variety with complex multiplication by a proper subfield of $\mathbb{Q}(\zeta_{p})$.
+  • $p$ does not divide $L\colon K]$ and $\operatorname{{\rm End}}^{0}(A)$ is isomorphic to a subfield of $\mathbb{Q}(\zeta_{p})$; or
+  • $p$ divides $L\colon K]$, $g\geq 3$ and $A$ is isogenous over $\overline{K}$ to the power of an absolutely simple abelian variety with complex multiplication by a proper subfield of $\mathbb{Q}(\zeta_{p})$.
 In particular there are only finitely many possibilities for $\operatorname{{\rm End}}^{0}(A)$.
@@
 Let $F/K$ be an odd degree abelian extension unramified outside of 2.
-Then by class field theory, any prime dividing $[F\colon K]$ divides either the class number of $K$, or the multiplicative order of the residue field of a prime above 2 in $K$.
+Then by class field theory, any prime dividing $F\colon K]$ divides either the class number of $K$, or the multiplicative order of the residue field of a prime above 2 in $K$.
 It follows that $K(A[2])/K$ is ramified at some prime $\mathfrak{q}\nmid 2$. This allows us to apply the above proposition and conclude. ∎
@@
 As $E/\mathbb{Q}$ is Galois, the reflex field $E^{*}$ is a subfield of $E$ [Shi98, Prop.
-28, pg 62].
+28, p.
+62].
 Moreover, since $A$ is absolutely simple its CM type is primitive.
-Proposition 30 on page 65 of [Shi98] applies to tell us the endomorphism field $L$ equals $E^{*}K$.
-Finally, if $E/\mathbb{Q}$ is an abelian extension, then $E=E^{*}$ by Example (1) on page 63 of [Shi98].
+Applying [Shi98, Prop.
+30, p.
+65] we find the endomorphism field $L$ equals $E^{*}K$.
+Finally, if $E/\mathbb{Q}$ is an abelian extension, then $E=E^{*}$ by [Shi98, Example (1), p.
+63].
 ∎
@@
   • $\operatorname{{\rm End}}^{0}(A)$ is a proper subfield of $\mathbb{Q}(\zeta_{p})$; or
-  • $p\in\{7,11,19,43,67,163\}$ and $A$ is isogenous over $\bar{\mathbb{Q}}$ to the power of an elliptic curve with complex multiplication by $\mathbb{Q}(\sqrt{-p})$.
+  • $p\in\{7,11,19,43,67,163\}$ and $A$ is isogenous over $\overline{\mathbb{Q}}$ to the power of an elliptic curve with complex multiplication by $\mathbb{Q}(\sqrt{-p})$.
 In particular there are only finitely many possibilities for $\operatorname{{\rm End}}^{0}(A)$.
@@
 Clearly, we may apply Theorem 2.4.
-Suppose first $p$ divides $[L:\mathbb{Q}]$.
-Then $g\geq 3$ and $A$ is isogenous over $\bar{\mathbb{Q}}$ to the power of an absolutely simple abelian variety with complex multiplication by a proper (non-trivial) subfield $F$ of $\mathbb{Q}(\zeta_{p})$.
+Suppose first $p$ divides $L\colon\mathbb{Q}$].
+Then $g\geq 3$ and $A$ is isogenous over $\overline{\mathbb{Q}}$ to the power of an absolutely simple abelian variety with complex multiplication by a proper (non-trivial) subfield $F$ of $\mathbb{Q}(\zeta_{p})$.
 Thus there is some minimal finite extension $K/\mathbb{Q}$ such that $A_{K}$ contains a simple abelian subvariety $B$ with CM by $F$.
 The above lemma applies to $B$ and shows $L$, the endomorphism field of $A$, contains $KF$.
-Proposition 3.9 of [Goo21] (recalled at the beginning of this section) shows $\operatorname{{\rm Gal}}(L/\mathbb{Q})$ is an extension of $C_{p}$ by $C_{2}^{m}$ for some $m$.
+By [Goo21, Prop.
+3.9] the group $\operatorname{{\rm Gal}}(L/\mathbb{Q})$ is an extension of $C_{p}$ by $C_{2}^{m}$ for some $m$.
 In particular, any non-trivial element of even order in $\operatorname{{\rm Gal}}(L/\mathbb{Q})$ has either order 2 or $2p$.
-Since $p$ does not divide $[F:\mathbb{Q}]$, the cyclic quotient $\operatorname{{\rm Gal}}(F/\mathbb{Q})$ of $\operatorname{{\rm Gal}}(L/\mathbb{Q})$ has order two.
+Since $p$ does not divide $F\colon\mathbb{Q}$], the cyclic quotient $\operatorname{{\rm Gal}}(F/\mathbb{Q})$ of $\operatorname{{\rm Gal}}(L/\mathbb{Q})$ has order two.
 Using $F$ is a CM field only ramified at $p$, we find $F=\mathbb{Q}(\sqrt{-p})$ where $p\equiv 3\mod{4}$.
 In particular $g$ is odd.
-We show $F$ has class number one, which by the Baker–Heegner–Stark Theorem will conclude the proof in this case.
+We shall now show $F$ has class number one, which by the Baker–Heegner–Stark Theorem will conclude the proof in this case.
 As $F/\mathbb{Q}$ is a quadratic extension, $B$ is an elliptic curve.
-Thus by CM theory [Sil94, II §4 Thm.
-4.1, pg 121] $KF$ contains $H$ the Hilbert class field of $F$.
-In particular, $L$ contains $H$, so $[H:F]$ divides $2^{m}p$ for some $m$.
-On the other hand, [FG18, Corollary 2.16] tells us every element of $\operatorname{{\rm Gal}}(H/F)$ has order dividing $g$.
+Thus by CM theory [Sil94, Thm.
+II.4.1, p.
+121] $KF$ contains $H$ the Hilbert class field of $F$.
+In particular, $L$ contains $H$, so $H\colon F]$ divides $2^{m}p$ for some $m$.
+On the other hand, [FG18, Cor.
+2.16] tells us every element of $\operatorname{{\rm Gal}}(H/F)$ has order dividing $g$.
 As $g$ is odd and less than $p$, we find $\operatorname{{\rm Gal}}(H/F)=1$ as claimed.
-We now suppose $p$ does not divide $[L:\mathbb{Q}]$ and show $E\neq\mathbb{Q}(\zeta_{p})$.
+We now suppose $p$ does not divide $L\colon\mathbb{Q}$] and show $E\neq\mathbb{Q}(\zeta_{p})$.
 Suppose we had equality, then applying Lemma 2.5 we find $L=\mathbb{Q}(\zeta_{p})$.
-If $p=3$, then we apply Theorem 2.9 of [Goo21] and Theorem 2.1 to obtain a contradiction.
-Else, $p\geq 5$ and Proposition 3.9 of [Goo21] provides us with a contradiction, since $\operatorname{{\rm Gal}}(\mathbb{Q}(\zeta_{p})/\mathbb{Q})$ is not an elementary abelian $2$-group.
+First let us note that as they are extensions of $\mathbb{Q}$ of coprime degree, we have $L\cap\mathbb{Q}(A[2])=\mathbb{Q}$.
+If $p=3$, then [Goo21, Thm.
+2.9] implies $\operatorname{{\rm End}}(A)$ is a 2-maximal order in $\operatorname{{\rm End}}^{0}(A)\cong\mathbb{Q}(\zeta_{3})$, so Theorem 2.1 gives us that $L\subseteq\mathbb{Q}(A[2])$.
+But $L\cap\mathbb{Q}(A[2])=\mathbb{Q}$ so $L=\mathbb{Q}$, contradicting the above.
+Thus we may take $p\geq 5$.
+Since $L\cap\mathbb{Q}(A[2])=\mathbb{Q}$, [Goo21, Prop.
+3.9] implies $\operatorname{{\rm Gal}}(L/\mathbb{Q})$ is an elementary abelian 2-group.
+However, $\operatorname{{\rm Gal}}(\mathbb{Q}(\zeta_{p})/\mathbb{Q})$ is not an elementary abelian $2$-group, thus we once again have a contradiction.
 ∎
-Let us record some easy corollaries of Theorem 2.9 of [Goo21], Theorems 2.1, 2.6 and Lemma 2.5.
+Let us record some easy corollaries of [Goo21, Thm. 2.9], Theorems 2.1, 2.6 and Lemma 2.5.
 Corollary 2.7.
-Let $A\colon y^{2}=f(x)$ be an elliptic curve defined over a number field with a real embedding. If $\operatorname{{\rm Gal}}(f)\cong C_{3}$, then $A$ does not have complex multiplication.
+Let $K$ be a number field with a real embedding.
+Let $f\in K[x]$ be an irreducible cubic polynomial, whose discriminant is a square in $K$.
+Then the elliptic curve defined by $y^{2}=f(x)$ does not have complex multiplication.
+Proof.
+Let $A$ denote the elliptic curve defined by $y^{2}=f(x)$.
+The assumptions on $f$ imply $\operatorname{{\rm Gal}}(f)\cong C_{3}$.
+Let us suppose for a contradiction that $A$ has complex multiplication.
+As $2$ is a primitive root modulo 3, we may apply [Goo21, Thm.
+2.9].
+Since $\mathbb{Q}(\zeta_{3})$ does not have any proper CM subfields, the second case of [Goo21, Thm.
+2.9] cannot hold.
+It follows that $(2)$ is inert in $E\coloneqq\operatorname{{\rm End}}^{0}(A)$ and $\operatorname{{\rm End}}(A)$ is 2-maximal in $E$.
+By Lemma 2.5, the endomorphism field of $A$ is $KE$.
+As $E$ does not have any real embeddings, it is not a subfield of $K$.
+Thus $KE\colon K]=2$.
+Since $K(A[2])=K(f)$ is an odd degree extension of $K$, it does not contain $KE$.
+Thus by Theorem 2.1 and the above, we find that $(2)$ is ramified in $E/\mathbb{Q}$.
+But this contradicts the above, so we find $A$ cannot have complex multiplication as claimed.
+∎
 Corollary 2.8.
@@
 Suppose $\operatorname{{\rm Gal}}(\mathbb{Q}(A[2])/\mathbb{Q})\cong C_{5}$.
-Then either $\operatorname{{\rm End}}(A)=\mathbb{Z}$ or $\operatorname{{\rm End}}_{\mathbb{Q}}^{0}(A)=\operatorname{{\rm End}}^{0}(A)=\mathbb{Q}(\sqrt{5})$.
+Then either $\operatorname{{\rm End}}(A)\cong\mathbb{Z}$ or $\operatorname{{\rm End}}_{\mathbb{Q}}^{0}(A)=\operatorname{{\rm End}}^{0}(A)\cong\mathbb{Q}(\sqrt{5})$.
+Proof.
+By Corollary 2.6, we have either $\operatorname{{\rm End}}^{0}(A)\cong\mathbb{Q}$ or $\mathbb{Q}(\sqrt{5})$.
+In the first case $\operatorname{{\rm End}}(A)\cong\mathbb{Z}$.
+Thus we may suppose $\operatorname{{\rm End}}^{0}(A)\cong\mathbb{Q}(\sqrt{5})$.
+We are left to show that $L$, the endomorphism field of $A$, equals $\mathbb{Q}$.
+The injection $\operatorname{{\rm Gal}}(L/\mathbb{Q})\hookrightarrow\operatorname{{\rm Aut}}_{\mathbb{Q}}(\operatorname{{\rm End}}^{0}(A))\cong\operatorname{{\rm Gal}}(\mathbb{Q}(\sqrt{5})/\mathbb{Q})$ implies the degree $L\colon\mathbb{Q}$] divides 2.
+Since $\mathbb{Q}(A[2])=\mathbb{Q}(f)$ is an odd degree extension of $\mathbb{Q}$, we find $L\cap\mathbb{Q}(A[2])=\mathbb{Q}$.
+Thus $L$ is contained in $\mathbb{Q}(A[2])$ if and only if $L=\mathbb{Q}$.
+As 2 is a primitive root modulo 5, we may apply [Goo21, Thm.
+2.9] which implies that $\operatorname{{\rm End}}(A)$ is 2-maximal in $\operatorname{{\rm End}}^{0}(A)\cong\mathbb{Q}(\sqrt{5})$.
+Since 2 is inert in $\mathbb{Q}(\sqrt{5})$, Theorem 2.1 implies that $L\subseteq\mathbb{Q}(A[2])$.
+Thus $L=\mathbb{Q}$ as claimed.
+∎
 Example 2.9.
@@
 Indeed, examples of the first case are easily found.
-For the second case, note the jacobian $J$ of the hyperelliptic curve $y^{2}=x(x^{5}-4x^{4}+2x^{3}+5x^{2}-2x-1)$ has $\operatorname{{\rm End}}_{\mathbb{Q}}(J)=\operatorname{{\rm End}}(J)\cong\mathbb{Z}\left[\frac{1+\sqrt{5}}{2}\right]$ by [Wil00, Proposition 1] and the Galois group of $x^{5}-4x^{4}+2x^{3}+5x^{2}-2x-1$ has order 5.
+For the second case, note the jacobian $J$ of the hyperelliptic curve $y^{2}=x(x^{5}-4x^{4}+2x^{3}+5x^{2}-2x-1)$ has $\operatorname{{\rm End}}_{\mathbb{Q}}(J)=\operatorname{{\rm End}}(J)\cong\mathbb{Z}\left[\frac{1+\sqrt{5}}{2}\right]$ by [Wil00, Prop.
+1] and the Galois group of $x^{5}-4x^{4}+2x^{3}+5x^{2}-2x-1$ has order 5.
 We present a variant of Theorem 2.4 for abelian varieties over imaginary quadratic fields.
@@
 Then either
-  • $\operatorname{{\rm End}}^{0}(A)$ is a proper subfield of $\mathbb{Q}(\zeta_{p})$; or
-  • $p\in\{7,11,19,43,67,163\}$ and $A$ is isogenous over $\bar{K}$ to the power of an elliptic curve with complex multiplication by $\mathbb{Q}(\sqrt{-p})$.
+  • $\operatorname{{\rm End}}^{0}(A)$ is isomorphic to a proper subfield of $\mathbb{Q}(\zeta_{p})$; or
+  • $p\in\{7,11,19,43,67,163\}$ and $A$ is isogenous over $\overline{K}$ to the power of an elliptic curve with complex multiplication by $\mathbb{Q}(\sqrt{-p})$.
 In particular there are only finitely many possibilities for $\operatorname{{\rm End}}^{0}(A)$.
@@
 Clearly, we may apply Theorem 2.4.
-Suppose first $p$ divides $[L:K]$.
-Then $A$ is isogenous over $\bar{K}$ to the power of an absolutely simple abelian variety $B$ with complex multiplication by a proper subfield $F$ of $\mathbb{Q}(\zeta_{p})$.
+Suppose first $p$ divides $L\colon K]$.
+Then $A$ is isogenous over $\overline{K}$ to the power of an absolutely simple abelian variety $B$ with complex multiplication by a proper subfield $F$ of $\mathbb{Q}(\zeta_{p})$.
 Lemma 2.5 applies to $B$ and shows $L$, the endomorphism field of $A$, contains $F$.
-Proposition 3.9 of [Goo21] (recalled at the beginning of this section) shows $\operatorname{{\rm Gal}}(L/K)$ is an extension of $C_{p}$ by $C_{2}^{m}$ for some $m$.
+By [Goo21, Prop.
+3.9] the group $\operatorname{{\rm Gal}}(L/K)$ is an extension of $C_{p}$ by $C_{2}^{m}$ for some $m$.
 In particular, any non-trivial element of even order in $\operatorname{{\rm Gal}}(L/K)$ has either order 2 or $2p$.
-Since $p$ does not divide $[F:\mathbb{Q}]$, the cyclic quotient $\operatorname{{\rm Gal}}(FK/K)$ of $\operatorname{{\rm Gal}}(L/K)$ has order dividing two.
-In turn we deduce $[F:\mathbb{Q}]$ divides $4$ and is equal to 4 only if $F\supsetneq K$.
-Being a CM field, $[F:\mathbb{Q}]=4$ would imply $F$ contains a real quadratic field.
-Whence $\operatorname{{\rm Gal}}(F/\mathbb{Q})\cong C_{2}\times C_{2}$ contradicting the fact $F\subseteq\mathbb{Q}(\zeta_{p})$.
-Thus $[F:\mathbb{Q}]=2$.
+Since $p$ does not divide $F\colon\mathbb{Q}$], the cyclic quotient $\operatorname{{\rm Gal}}(FK/K)$ of $\operatorname{{\rm Gal}}(L/K)$ has order dividing two.
+In turn we deduce $F\colon\mathbb{Q}$] divides $4$ and is equal to 4 only if $F\supsetneq K$.
+Being a CM field, $F\colon\mathbb{Q}]=4$ would imply $F$ contains a real quadratic field in addition to the imaginary quadratic field $K$.
+Thus $\operatorname{{\rm Gal}}(F/\mathbb{Q})\cong C_{2}\times C_{2}$ contradicting the fact $F\subseteq\mathbb{Q}(\zeta_{p})$.
+Thus $F\colon\mathbb{Q}]=2$.
 As $F$ is a CM field only ramified at $p$, we deduce $F=\mathbb{Q}(\sqrt{-p})$ where $p\equiv 3\mod{4}$.
-[…]
+Let $H$ be the Hilbert class field of $F=\mathbb{Q}(\sqrt{-p})$.
+By [FG18, Thm.
+2.14] the group $\operatorname{{\rm Gal}}(HK/FK)$ has order dividing $g$, thus arguing as in the proof of Theorem 2.6, we deduce $HK\colon FK]=1$.
+In other words, $H\subseteq FK=K(\sqrt{-p})$.
 Thus either $H=\mathbb{Q}(\sqrt{-p})$ or $H=K(\sqrt{-p})$.
@@
 As $p\equiv 3\mod{4}$ and $g\geq 2$, it follows, from the classification of imaginary quadratic fields of class number one, that $p\in\{7,11,19,43,67,163\}$.
-We now suppose $p$ does not divide $[L:K]$ and show $E\neq\mathbb{Q}(\zeta_{p})$.
+We now suppose $p$ does not divide $L\colon K]$ and show $E\neq\mathbb{Q}(\zeta_{p})$.
 Suppose we had equality, then by Lemma 2.5, $L=K(\zeta_{p})$.
-Proposition 3.9 of [Goo21] then implies $\operatorname{{\rm Gal}}(L/K)=\operatorname{{\rm Gal}}(K(\zeta_{p})/K)\cong\operatorname{{\rm Gal}}(\mathbb{Q}(\zeta_{p})/\mathbb{Q}(\zeta_{p})\cap K)$ has order at most $2$.
-Hence $[\mathbb{Q}(\zeta_{p}):\mathbb{Q}]$ divides $4$, being equal to 4 only if $K$ is contained in $\mathbb{Q}(\zeta_{p})$.
-But $[\mathbb{Q}(\zeta_{p}):\mathbb{Q}]=4$ implies $p=5$ and $\mathbb{Q}(\zeta_{5})$ does not have an imaginary subfield.
-Whence $[\mathbb{Q}(\zeta_{p}):\mathbb{Q}]=2$ and $p=3$.
+Applying [Goo21, Prop.
+3.9] shows $\operatorname{{\rm Gal}}(L/K)=\operatorname{{\rm Gal}}(K(\zeta_{p})/K)\cong\operatorname{{\rm Gal}}(\mathbb{Q}(\zeta_{p})/\mathbb{Q}(\zeta_{p})\cap K)$ has order at most $2$.
+Hence $\mathbb{Q}(\zeta_{p})\colon\mathbb{Q}$] divides $4$, being equal to 4 only if $K$ is contained in $\mathbb{Q}(\zeta_{p})$.
+But $\mathbb{Q}(\zeta_{p})\colon\mathbb{Q}]=4$ implies $p=5$ and $\mathbb{Q}(\zeta_{5})$ does not have an imaginary subfield.
+Whence $\mathbb{Q}(\zeta_{p})\colon\mathbb{Q}]=2$ and $p=3$.
 This in turn implies $g=1$ which we have ruled out by assumption.
@@
 Question 2.12.
-Let $K$ be an imaginary quadratic field of class number divisible by 5.
+Let $K$ be an imaginary quadratic field with class number divisible by 5.
 Let $H$ be a degree 5 extension of $K$ contained in the Hilbert class field of $K$.
@@
 Remark 2.13.
-By combining Theorem 2.1, Theorem 2.10 of [Goo21] and [Shi98, Prop.
-30, pg.65; Example (2), pg.64], we see that either $\operatorname{{\rm End}}^{0}(A)\cong\mathbb{Q}$ or $\mathbb{Q}(\sqrt{d})$ with $d\equiv 5\mod{8}$.
+With reference to the above question, there are already several restrictions on the possible endomorphism algebras.
+Indeed, by combining Theorem 2.1, [Goo21, Thm.
+2.10] and [Shi98, Prop.
+30, p.
+65; Example (2), p.
+64], we see that either $\operatorname{{\rm End}}^{0}(A)\cong\mathbb{Q}$ or $\mathbb{Q}(\sqrt{d})$ with $d\equiv 5\mod{8}$.
 The above question then becomes: given $H$, which $d$ occur?
-We finish this section with the following extension of [Goo21, Theorem 3.5].
+We finish this section with the following extension of [Goo21, Thm. 3.5].
 Theorem 2.14.
-Let $f\in K[x]$ be a polynomial of odd degree $n$ with Galois group isomorphic to a Frobenius group $G$ of order $n(n-1)$.
+Let $f\in K[x]$ be a polynomial of odd degree $n$ with Galois group isomorphic to a Frobenius group $G$ of order n(n1)n(n-1).
 Let $J_{f}$ be the jacobian associated to the hyperelliptic curve defined by $y^{2}=f(x)$.
-Suppose $E={\rm End}^{0}(J_{f})$ is a number field of dimension $s$ over $\mathbb{Q}$.
+Suppose ${\rm End}^{0}(J_{f})$ is isomorphic to a number field $E$ of dimension $s$ over $\mathbb{Q}$.
 Then $E/\mathbb{Q}$ is Galois with ${\rm Gal}(E/\mathbb{Q})$ isomorphic to a quotient of $H$ the Frobenius complement of $G$.
-Furthermore, $L/K$ is an extension of degree $s$ contained in $K(f)$, and as abstract groups, ${\rm Gal}(L/K)\cong{\rm Gal}(E/\mathbb{Q})$. Moreover, if $s=n-1$, then $L=EK$.
+Furthermore, $L/K$ is an extension of degree $s$ contained in $K(f)$, and as abstract groups, ${\rm Gal}(L/K)\cong{\rm Gal}(E/\mathbb{Q})$. Moreover, if s=n1s=n-1, then $L=EK$.
 Finally, if $\operatorname{{\rm End}}(J_{f})$ is $2$-maximal, then $E$ is unramified at 2.
 Proof.
-This result is [Goo21, Theorem 3.5] with the extra assertion that if $s=n-1$, then $L=EK$.
-If $s=n-1$, then $J_{f}$ is an absolutely simple abelian variety with CM by $E$.
+This result is [Goo21, Thm.
+3.5] with the extra assertion that if s=n1s=n-1, then $L=EK$.
+If s=n1s=n-1, then $J_{f}$ is an absolutely simple abelian variety with CM by $E$.
 We may therefore apply Lemma 2.5 to find $L=E^{*}K$ and $E^{*}\subseteq E$.
-Hence $[E^{*}:\mathbb{Q}]\geq[L:K]=s=[E:\mathbb{Q}]$.
+Hence $E^{*}\colon\mathbb{Q}]\geq[L\colon K]=s=[E\colon\mathbb{Q}$].
 It follows we have equality $E^{*}=E$.
@@
 The unique degree 4 extension contained in $\mathbb{Q}(f)$ is given by $K$.
-This field is totally inert at 2 (as predicted by [Goo21, Theorem 2.9]) and unramified outside 13.
+This field is totally inert at 2 (as predicted by [Goo21, Thm.
+2.9]) and unramified outside 13.
 We note its class number is one, but in line with the theorems presented above, $\mathbb{Q}(f)/K$ is a degree 5 extension ramified only at 2.
 3. An analogue of Theorem 1.2 for QM surfaces
-Let $B=\mathbb{Q}+i\mathbb{Q}+j\mathbb{Q}+k\mathbb{Q}$ be an indefinite quaternion algebra generated by $i,j$ satisfying $i^{2}=D/m$, $j^{2}=m$, $ij=-ji$ and $k=ij$, where $D$ is a positive squarefree integer and $m|D$.
+Let $B=\mathbb{Q}+i\mathbb{Q}+j\mathbb{Q}+k\mathbb{Q}$ be an indefinite quaternion algebra generated by $i,j$ satisfying $i^{2}=D/m$, $j^{2}=m$, ij=jiij=-ji and $k=ij$, where $D$ is a positive squarefree integer and $m|D$.
 We note that as $B$ is indefinite, $0>-m=\mathrm{n}(j)$, the reduced norm of $j$.
 As we shall only use the reduced norm, reduced trace and reduced discriminant, we shall refer to these simply as the norm, trace and discriminant respectively.
-An order of $B$ is said to be hereditary if it has squarefree (reduced) discriminant (for alternate, equivalent, definitions see [DR04, Voi21]).
+An order of $B$ is said to be hereditary if it has squarefree discriminant (for alternate, equivalent, definitions see [DR04, Voi21]).
 We shall take $\mathcal{O}$ to be a hereditary order in $B$ of discriminant $D$.
@@
 We shall assume there is an element $\mu\in\mathcal{O}$ satisfying $\mu^{2}+D=0$.
-This is always the case when $\mathcal{O}$ is a maximal order [Voi21, 43.6.6, page 818] and furthermore such an element $\mu$ induces a principal polarisation on $A$ [Voi21, page 821].
+This is always the case when $\mathcal{O}$ is a maximal order [Voi21, 43.6.6, p.
+818] and furthermore such an element $\mu$ induces a principal polarisation on $A$ [Voi21, p.
+821].
 Fix such an $A$, $\mu$ and $\iota$.
 Following [DR04, Definition 3.3], we say $\chi\in B$ is a twist of $(\mathcal{O},\mu)$ if it lies in both $\mathcal{O}$ and the normaliser $N_{B^{*}}(\mathcal{O})$, has trace zero, $\mathrm{n}(\chi)$ divides $D$, and $\mu\chi=-\chi\mu$.
-The existence of such an element can be verified by a finite computation, we refer the reader to the remark at the bottom of page 9 of [DR04] for further details.
+The existence of such an element can be verified by a finite computation, we refer the reader to [DR04, Rem., p.
+9] for further details.
 Owing in part to the fact $(\mathcal{O},\mu)$ has been fixed, there are only two possible values [^2] This follows from [Rot04, Lemmas 3.5 and 3.7].
@@
 This allows us to write $V_{0}=\langle[\mu],[\chi]\rangle$ where the representative $\chi$ may be taken to have reduced norm $m|D$.
-Moreover, any representative of $[\chi]$ has reduced norm $m$ up to a rational square.
-Likewise, representatives of the class $[\mu\chi]$ have reduced norm $D/m$ up to rational squares.
+Moreover, any representative of $\chi]$ has reduced norm $m$ up to a rational square.
+Likewise, representatives of the class $\mu\chi]$ have reduced norm $D/m$ up to rational squares.
 $\mathrm{n}(\chi)$ can take and the product of these values is equal to $D$.
-We call $(\mathcal{O},\mu,\chi)$ a twisted principally polarised order, and say $(\mathcal{O},\mu,\chi)$ is of discriminant $D$ and norm $-\mathrm{n}(\chi)$.
+We call $(\mathcal{O},\mu,\chi)$ a twisted principally polarised order, and say $(\mathcal{O},\mu,\chi)$ is of discriminant $D$ and norm $\mathrm{n}(\chi)$.
 As before, we let $L$ denote the endomorphism field of $A$.
-Dieulefait and Rotger [DR04, Theorem 3.4] showed $\operatorname{{\rm Gal}}(L/K)$ is isomorphic to one of the trivial group, $C_{2}$, or $C_{2}\times C_{2}$.
+Dieulefait and Rotger [DR04, Thm.
+3.4] showed $\operatorname{{\rm Gal}}(L/K)$ is isomorphic to one of the trivial group, $C_{2}$, or $C_{2}\times C_{2}$.
 Moreover, they proved in each case $\operatorname{{\rm End}}_{K}^{0}(A)$ is respectively isomorphic to $B$; one of $\mathbb{Q}(\mu),\mathbb{Q}(\chi),\mathbb{Q}(\mu\chi)$; or $\mathbb{Q}$.
 This determines (and heavily restricts) the possible images of
-$\operatorname{{\rm Gal}}(L/K)\rightarrow\operatorname{{\rm Aut}}(\operatorname{{\rm End}}(A))\cong N_{B^{*}}(\mathcal{O})/\mathbb{Q}^{*}$
+$\operatorname{{\rm Gal}}(L/K)\hookrightarrow\operatorname{{\rm Aut}}_{\mathbb{Z}}(\operatorname{{\rm End}}(A))\cong N_{B^{*}}(\mathcal{O})/\mathbb{Q}^{*}$
 arising from the natural action of $\operatorname{{\rm Gal}}(L/K)$ on $\operatorname{{\rm End}}(A)$.
@@
 is an order of discriminant $D$ in $B$.
-Moreover, any order $\mathcal{O}^{\prime}$ of discriminant $D$ which contains $\mathbb{Z}[1,mi,j,k]$, satisfies $\mathcal{O}^{\prime}\otimes\mathbb{Z}_{2}=\mathcal{O}\otimes\mathbb{Z}_{2}$.
+Moreover, any order $\mathcal{O}^{\prime}$ of discriminant $D$ which contains $\mathbb{Z}[1,mi,j,k]$, satisfies $\mathcal{O}^{\prime}\otimes\mathbb{Z}_{(2)}=\mathcal{O}\otimes\mathbb{Z}_{(2)}$.
 Proof.
-It is a routine calculation to show $\mathcal{O}$ is an order of discriminant $D$ (for an example, see [Vig80, pg. 85 - 86]).
-As $m$ is odd we have $\mathbb{Z}[1,mi,j,k]\otimes\mathbb{Z}_{2}=\mathbb{Z}[1,i,j,k]\otimes\mathbb{Z}_{2}$.
-Thus we may assume $\mathcal{O}^{\prime}$ contains $\mathbb{Z}[1,i,j,k]$.
-The discriminant of $\mathbb{Z}[1,i,j,k]$ equals $4D$ and thus any element contained in $\mathcal{O}^{\prime}$ but not in $\mathbb{Z}[1,i,j,k]$ is of the form $\frac{1}{4}\alpha$ with $\alpha\in\mathbb{Z}[1,i,j,k]$.
+It is a routine calculation to show $\mathcal{O}$ is an order of discriminant $D$ (for an example, see [Vig80, p. 85 - 86]).
+As $m$ is odd we have $\mathbb{Z}[1,i,j,k]\otimes\mathbb{Z}_{(2)}=\mathbb{Z}[1,mi,j,k]\otimes\mathbb{Z}_{(2)}\subseteq\mathcal{O}^{\prime}\otimes\mathbb{Z}_{(2)}$.
+There exists a lattice $\mathcal{O}^{\prime\prime}$ such that $\mathcal{O}^{\prime\prime}\otimes\mathbb{Z}_{(2)}=\mathcal{O}^{\prime}\otimes\mathbb{Z}_{(2)}$ and $\mathcal{O}^{\prime\prime}\otimes\mathbb{Z}_{(p)}=\mathbb{Z}[1,i,j,k]\otimes\mathbb{Z}_{(p)}$ for every odd prime $p$ [Voi21, Thm.
+9.4.9].
+Moreover, $\mathcal{O}^{\prime\prime}$ is in fact an order since every localisation is [Voi21, Lem.
+10.2.10].
+As $\mathbb{Z}[1,i,j,k]\otimes\mathbb{Z}_{(p)}$ is contained in $\mathcal{O}^{\prime\prime}\otimes\mathbb{Z}_{(p)}$ for every prime $p$, we deduce $\mathbb{Z}[1,i,j,k]$ is contained in $\mathcal{O}^{\prime\prime}$ [Voi21, Cor.
+9.4.7].
+The discriminant of $\mathbb{Z}[1,i,j,k]$ equals $4D$.
+As the discriminant of an order is determined locally [Voi21, 15.2.13], we deduce the discriminant of $\mathcal{O}^{\prime\prime}$ equals $D$.
+Applying [Voi21, Lem.
+15.2.5], we see that any element contained in $\mathcal{O}^{\prime\prime}$ but not in $\mathbb{Z}[1,i,j,k]$ is of the form $\frac{1}{4}\alpha$ with $\alpha\in\mathbb{Z}[1,i,j,k]$.
 Let us write such an element as $\frac{1}{4}\alpha=w+xi+yj+zk$.
@@
 We have $\mathrm{n}(a+bi+cj+dk)\equiv a-2b+c+2d\equiv 0\mod{4}$, from which we deduce $\frac{1}{2}(1+j+k)$, $\frac{1}{2}(1+i+j)$, $\frac{1}{2}(i+k)$ are the only integral such.
-In order for the discriminant of $\mathcal{O}^{\prime}$ to equal $D$ we see that all of these elements must belong to $\mathcal{O}^{\prime}$.
+In order for the discriminant of $\mathcal{O}^{\prime\prime}$ to equal $D$ we see that all of these elements must belong to $\mathcal{O}^{\prime\prime}$.
+Whence $\mathcal{O}\otimes\mathbb{Z}_{(2)}=\mathcal{O}^{\prime\prime}\otimes\mathbb{Z}_{(2)}=\mathcal{O}^{\prime}\otimes\mathbb{Z}_{(2)}$.
 ∎
@@
 is an order of discriminant $D$ in $B$.
-Moreover, any order $\mathcal{O}^{\prime}$ of even discriminant $D$ which contains $\mathbb{Z}[1,mi,j,k]$ satisfies $\mathcal{O}^{\prime}\otimes\mathbb{Z}_{2}=\mathcal{O}\otimes\mathbb{Z}_{2}$.
-Any order $\mathcal{O}^{\prime}$ of discriminant $D\equiv t\mod{4}$ with $t\in\{1,3\}$, which contains $\mathbb{Z}[1,mi,j,k]$, satisfies either $\mathcal{O}^{\prime}\otimes\mathbb{Z}_{2}=\mathcal{O}\otimes\mathbb{Z}_{2}$ or $\mathcal{O}^{\prime}\otimes\mathbb{Z}_{2}=\mathcal{O}_{t}\otimes\mathbb{Z}_{2}$.
+Moreover, any order $\mathcal{O}^{\prime}$ of even discriminant $D$ which contains $\mathbb{Z}[1,mi,j,k]$ satisfies $\mathcal{O}^{\prime}\otimes\mathbb{Z}_{(2)}=\mathcal{O}\otimes\mathbb{Z}_{(2)}$.
+Any order $\mathcal{O}^{\prime}$ of discriminant $D\equiv t\mod{4}$ with $t\in\{1,3\}$, which contains $\mathbb{Z}[1,mi,j,k]$, satisfies either $\mathcal{O}^{\prime}\otimes\mathbb{Z}_{(2)}=\mathcal{O}\otimes\mathbb{Z}_{(2)}$ or $\mathcal{O}^{\prime}\otimes\mathbb{Z}_{(2)}=\mathcal{O}_{t}\otimes\mathbb{Z}_{(2)}$.
 Proof.
@@
 Proof.
-As $\operatorname{{\rm End}}(A)$ is a hereditary order, the results of [DR04, Theorem 3.4] apply to show a non-trivial element of $\operatorname{{\rm Gal}}(L/K)$ acts on $\operatorname{{\rm End}}(A)$ (possibly after scaling) by conjugation as one of $\mu,\chi$ or $\mu\chi$.
+If $L=K$, then there is nothing left to show.
+Thus let us suppose $L\neq K$.
+As $\operatorname{{\rm End}}(A)$ is a hereditary order, the results of [DR04, Thm.
+3.4] apply to show $\operatorname{{\rm Gal}}(L/K)\leq C_{2}^{2}$ and a non-trivial element of $\operatorname{{\rm Gal}}(L/K)$ acts on $\operatorname{{\rm End}}(A)$ (possibly after scaling) by conjugation as one of $\mu,\chi$ or $\mu\chi$.
 We look to determine the action of these elements on $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}/2\mathbb{Z}$.
-To do so we reduce a $\mathbb{Z}$-basis of $\operatorname{{\rm End}}(A)$ modulo $2$, thus we may work with $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}_{2}$ in place of $\operatorname{{\rm End}}(A)$.
+To do so we shall consider a $\mathbb{Z}_{(2)}$-basis of $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}_{(2)}$ reduced modulo $2$.
 By considering the algebraic relations they satisfy, we may assume $\mu=k$, $\chi=j$ and $\mu\chi=mi$.
-Hence Lemma 3.1 allows us to take $\operatorname{{\rm End}}(A)=\mathbb{Z}+\frac{1}{2}(1+j+k)\mathbb{Z}+\frac{1}{2}(1+j-k)\mathbb{Z}+\frac{1}{2}(i+k)\mathbb{Z}$.
+Hence Lemma 3.1 allows us to take $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}_{(2)}=\mathbb{Z}_{(2)}+\frac{1}{2}(1+j+k)\mathbb{Z}_{(2)}+\frac{1}{2}(1+j-k)\mathbb{Z}_{(2)}+\frac{1}{2}(i+k)\mathbb{Z}_{(2)}$.
 Let $X=\frac{1}{2}(1+j+k)$, $Y=\frac{1}{2}(1+j-k)$ and $Z=\frac{1}{2}(i+k)$.
-Let us examine the action of $i,j$ and $k$ on the basis of $\operatorname{{\rm End}}(A)$ given by $1,X,Y,Z$.
+Let us examine the action of $i,j$ and $k$ on the $\mathbb{Z}_{(2)}$-basis of $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}_{(2)}$ given by $1,X,Y,Z$.
 Each of $i,j$ and $k$ fix $1$.
@@
 Looking at the coefficients, we see the action remains faithful on $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}/2\mathbb{Z}$.
-By proceeding as in the proof of Theorem 2.1 we are now done.
+Recall that the natural map $\operatorname{{\rm End}}(A)\rightarrow\operatorname{{\rm End}}(A[2])$ induces an injection $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}/2\mathbb{Z}\hookrightarrow\operatorname{{\rm End}}(A[2])$.
+The group $\operatorname{{\rm Gal}}(\overline{K}/K(A[2]))$ acts trivially on $\operatorname{{\rm End}}(A[2])$ and hence trivially on $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}/2\mathbb{Z}$ as well.
+Since $\operatorname{{\rm Gal}}(L/K)$ acts faithfully on $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}/2\mathbb{Z}$, we find $L\subseteq K(A[2])$.
 ∎
@@
 Proof.
-As $\operatorname{{\rm End}}(A)$ is a hereditary order, the results of [DR04, Theorem 3.4] apply to show a non-trivial element of $\operatorname{{\rm Gal}}(L/K)$ acts on $\operatorname{{\rm End}}(A)$ (possibly after scaling) by conjugation as one of $\mu,\chi$ or $\mu\chi$.
+If $L=K$, then there is nothing left to show.
+Thus let us suppose $L\neq K$.
+As $\operatorname{{\rm End}}(A)$ is a hereditary order, the results of [DR04, Thm.
+3.4] apply to show $\operatorname{{\rm Gal}}(L/K)\leq C_{2}^{2}$ and a non-trivial element of $\operatorname{{\rm Gal}}(L/K)$ acts on $\operatorname{{\rm End}}(A)$ (possibly after scaling) by conjugation as one of $\mu,\chi$ or $\mu\chi$.
 We look to determine the action of these elements on $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}/2\mathbb{Z}$.
-To do so we reduce a $\mathbb{Z}$-basis of $\operatorname{{\rm End}}(A)$ modulo $2$, thus we may work with $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}_{2}$ in place of $\operatorname{{\rm End}}(A)$.
+To do so we shall consider a $\mathbb{Z}_{(2)}$-basis of $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}_{(2)}$ reduced modulo $2$.
 By considering the algebraic relations they satisfy, we may assume $\mu=k$, $\chi=j$ and $\mu\chi=mi$.
-Lemma 3.2 allows us to take $\operatorname{{\rm End}}(A)$ equal to one of the three given orders.
-As these orders differ by permuting $i,j$ and $k$, we only give details for the case $\operatorname{{\rm End}}(A)=\mathbb{Z}+\frac{1}{2}(1+j)\mathbb{Z}+k\mathbb{Z}+\frac{1}{2}(i+k)\mathbb{Z}$.
+Lemma 3.2 allows us to take $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}_{(2)}$ equal to one of the three given orders tensored by $\mathbb{Z}_{(2)}$.
+As these orders differ by permuting $i,j$ and $k$, we only give details for the case $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}_{(2)}=\mathbb{Z}_{(2)}+\frac{1}{2}(1+j)\mathbb{Z}_{(2)}+k\mathbb{Z}_{(2)}+\frac{1}{2}(i+k)\mathbb{Z}_{(2)}$.
 Let $X=\frac{1}{2}(1+j)$, $Y=k$ and $Z=\frac{1}{2}(i+k)$.
-Let us examine the action of $i,j$ and $k$ on the basis of $\operatorname{{\rm End}}(A)$ given by $1,X,Y,Z$.
+Let us examine the action of $i,j$ and $k$ on the $\mathbb{Z}_{(2)}$-basis of $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}_{(2)}$ given by $1,X,Y,Z$.
 Each of $i,j$ and $k$ fix $1$.
@@
 Looking at the coefficients, we see $j$ acts trivially on $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}/2\mathbb{Z}$ whereas $i$ and $k$ act by the same involution.
-Proceeding as in the proof of Theorem 2.1, we find at least one of $\mathbb{Q}(i)\cong\mathbb{Q}(\sqrt{D/m})$ and $\mathbb{Q}(k)\cong\mathbb{Q}(\sqrt{-D})$ is contained in $\operatorname{{\rm End}}_{F}^{0}(A)$.
+We thus find that $\operatorname{{\rm Gal}}(L/K)$ acts on $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}/2\mathbb{Z}$ through a quotient group $\operatorname{{\rm Gal}}(L^{\prime}/K)$ of order at most 2.
+As $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}/2\mathbb{Z}\hookrightarrow\operatorname{{\rm End}}(A[2])$, we find $L^{\prime}\subseteq K(A[2])$.
+The above calculation shows the action of $\operatorname{{\rm Gal}}(L^{\prime}/K)$ is via conjugation by one of either $i$ or $k$.
+Hence, we find at least one of $\mathbb{Q}(i)\cong\mathbb{Q}(\sqrt{D/m})$ and $\mathbb{Q}(k)\cong\mathbb{Q}(\sqrt{-D})$ is contained in $\operatorname{{\rm End}}_{F}^{0}(A)$, see the proof of [DR04, Thm.
+3.4].
 ∎